### v1

# From Evidence to Verdict: An Agent-Based Forensic Framework for AI-Generated Image Detection

Abstract
The rapid evolution of AI-generated images poses unprecedented challenges to information integrity and media authenticity. Existing detection approaches suffer from fundamental limitations: traditional classifiers lack interpretability and fail to generalize across evolving generative models, while vision-language models (VLMs), despite their promise, remain constrained to single-shot analysis and pixel-level reasoning. To address these challenges, we introduce AIFo (Agent-based Image Forensics), a novel training-free framework that emulates human forensic investigation through multi-agent collaboration. Unlike conventional methods, our framework employs a set of forensic tools, including reverse image search, metadata extraction, pre-trained classifiers, and VLM analysis, coordinated by specialized LLM-based agents that collect, synthesize, and reason over cross-source evidence. When evidence is conflicting or insufficient, a structured multi-agent debate mechanism allows agents to exchange arguments and reach a reliable conclusion. Furthermore, we enhance the framework with a memory-augmented reasoning module that learns from historical cases to improve future detection accuracy. Our comprehensive evaluation spans 6,000 images across both controlled laboratory settings and challenging real-world scenarios, including images from modern generative platforms and diverse online sources. AIFo achieves 97.05% accuracy, substantially outperforming traditional classifiers and state-of-the-art VLMs. These results demonstrate that agent-based procedural reasoning offers a new paradigm for more robust, interpretable, and adaptable AI-generated image detection.
Introduction
In recent years, image generative models such as GLIDE [37], Imagen [45], DALL·E 2 [43], and Stable Diffusion [44], have advanced rapidly. They can synthesize photorealistic images from natural language in seconds [23, 44, 45]. However, the realism of AI-generated images has raised serious societal concerns. Because AI-generated images can now easily fool human observers, malicious actors are increasingly leveraging them to spread disinformation and impersonate individuals [48]. For example, during the 2024 U.S. presidential election cycle, sophisticated deepfakes (i.e., AI-generated images or videos that convincingly fabricate real people or events) have appeared in campaign ads and on social media, which potentially manipulate public opinions and disrupt the voting behaviors [15]. Beyond elections, AI-generated images also introduce broader risks [48, 50], including misinformation and privacy infringement.
In response to these risks, substantial research has been devoted to the detection of AI-generated images. Current methodologies can be generally classified into two main categories: traditional machine learning classifiers and advanced approaches leveraging large vision language models (VLMs).
Traditional machine learning classifiers typically rely on training convolutional neural networks (CNNs) or transformer-based models to distinguish between real and fake images [53, 59, 54, 19, 17]. Early studies reveal that AI-generated images tend to exhibit shared low-level artifacts, allowing detectors trained on labeled images to identify them [53, 46]. For example, DE-FAKE [46] trains a set of classifiers on AI-generated and real images to learn AI-specific artifacts. However, these works often depend on a limited number of training datasets with fake images produced by only a few specific generative models, which makes them prone to generalizability issues: they perform well on seen data but struggle to generalize to unseen images from new generative models [60, 51]. They also often lack explainability: most models act as black boxes, producing binary outputs without offering human-interpretable justifications [33, 60].
More recently, vision language models (VLMs) have shown promise for more generalizable and explainable detection [60, 57, 33, 28]. Due to the large-scale pre-training, VLMs can be transferred to image detection tasks in a zero-shot or few-shot manner [60, 57, 51], without relying on specialized labeled datasets. Beyond identifying pixel-level artifacts like traditional classifiers, VLMs can also apply semantic and world knowledge in the detection process, e.g., a photo of a flying cat is AI-generated because a flying cat cannot exist in the physical world. In addition, they can provide human-interpretable justifications through prompt engineering for explainable detection [57, 28, 51].
Despite these advancements, both traditional and VLM-based approaches suffer from fundamental limitations compared to human forensic experts. First, they rely heavily on pixel-level image features as the main detection evidence. By contrast, human experts can not only employ off-the-shelf classifiers and VLMs for pixel-level analysis, but also actively seek evidence beyond the images themselves, e.g., Exchangeable Image File Format (EXIF) metadata that contains important information about the image or online contextual information to make a comprehensive decision. Second, both approaches treat image detection as a single-shot classification task, using a fixed model or pipeline. Human experts, on the other hand, approach image detection as a dynamic reasoning process: they flexibly employ different tools and iteratively refine their judgments based on the gathered evidence. Finally, human experts can improve over time through accumulated experience, enabling them to adapt to new generative models and real-world scenarios, whereas fixed models remain static unless frequently fine-tuned on new datasets. Motivated by this, we aim to explore a new paradigm that combines the advantages of existing classifiers and VLMs with the capabilities of human experts for AI-generated image detection.
Our Work. Our approach fundamentally differs from conventional methods based on static classifiers or single VLM. Rather than developing another standalone image classifier, we design a cognitive AI agent system that automates the entire human-like forensic workflow, simulating reasoning and decision-making in the detection of AI-generated images. We present AIFo (Agent-based Image Forensics), an LLM-based multi-agent framework that integrates functionalities such as visual recognition, semantic understanding, and provenance analysis. AIFo consists of a forensic Toolbox, an Evidence Gatherer, a Reasoning Agent, and a Debate module. Given a test image, the Evidence Gatherer first applies tools from the Toolbox and aggregates the resulting evidence. The Reasoning Agent then evaluates the quality of the collected evidence, i.e., whether it is sufficient and consistent enough to support a reliable judgment. If so, the Reasoning Agent produces a final decision and explains why the image is classified as AI-generated or real. Otherwise, we introduce a Debate module [21, 35, 36, 34, 49] to handle cases where the evidence is insufficient or conflicting. In this module, two Debate Agents exchange arguments over multiple rounds, each adopting an opposing stance (for or against the claim that the image is AI-generated), while a Judge Agent oversees the debate process and produces the final judgment. Additionally, we demonstrate the potential of our framework to continuously and progressively enhance its detection effectiveness by learning from historical testing data. By incorporating a memory module that stores all testing history, the framework behaves like a human expert, accumulating experience and improving performance over time. Overall, such a cognitive AI agent framework offers a new paradigm for AI-generated image detection.
Main Findings. We evaluate AIFo on a dataset of 6,000 AI-generated and real images, comprising 3,000 samples from five established benchmarks (e.g., Flickr30K [40], GenImage [61], and FakeBench [33]) and 3,000 in-the-wild images collected from six online platforms. AIFo is benchmarked against a range of baselines methods, including traditional classifiers such as CNNSpot [53], DE-FAKE [46], and state-of-the-art VLMs (e.g., GPT-4.1 [8] and GPT-4o [39]). To ensure a fair comparison, we disable the memory module and prevent our framework from learning during the main evaluation. Even under these conditions, AIFo achieves the best overall performance, reaching 0.9705 accuracy and surpassing GPT-4o (0.9483), GPT-4.1 (0.9416), and other baselines. It also maintains higher robustness (0.9047–0.9690) than GPT-4o (0.8818–0.9462) under three types of perturbations. To demonstrate the effectiveness of the memory module, we conduct a case study on 50 images that are misclassified in the main evaluation. The results show that when the memory module stores similar historical cases, approximately 40% of these errors are successfully corrected.
The main contributions of our work are as follows:
  • Human-Like Procedural Reasoning. We present the first agentic framework (AIFo) that simulates human-like procedural reasoning for AI-generated image detection. Our framework combines the strengths of conventional classifiers and VLMs, employing them as complementary tools, and incorporates human-like reasoning processes such as integrating multiple sources of evidence, debating conflicting evidence, and improving through accumulated experience. This highlights a transition from static classification toward a dynamic reasoning process in AI-generated image detection.
  • Comprehensive Evaluation. We contribute a valuable benchmark dataset covering 6000 images from both existing benchmark datasets and in-the-wild images from internet platforms. This dataset serves as the foundation for rigorous evaluation, through which we demonstrate the superior performance, generalizability, and robustness of our AIFo compared to state-of-the-art baselines.
  • Training-Free Core and Cross-Model Generalizability. Unlike conventional detectors that require large-scale training data and frequent updates, the core of AIFo is training-free and designed to generalize across evolving generative models. It leverages generalizable forensic evidence rather than model-specific features, enabling robust detection of AI-generated images. An optional memory module further enhances performance over time by accumulating historical cases.
Threat Model
Detector’s Goals
Distinguishing Between AI-Generated and Real Images. The detector aims to determine whether a given image is AI-generated or a real image. In this work, AI-generated images include those produced by generative models like text-to-image diffusion models, as well as AI-edited images, such as cartoons generated from real photographs and images replaced with AI-generated backgrounds. Real images refer to those captured by physical cameras or manually created/drawn by human artists. Notably, if a real image has been edited or post-processed by humans (without the aid of AI models), e.g., by applying filters, adjusting color contrast, or increasing brightness, we still consider it a real image.
Enhanced Interpretability. Unlike conventional binary detectors, our detector aims not only to provide a decision but also to offer human-interpretable justifications. For example, if our detector identifies an image as an AI-generated image, it should also point out evidence that supports this decision, such as unreasonable object placement (semantic-level evidence), distorted lighting in the image (pixel-level evidence), editing history (metadata-level evidence), AI watermarks (source-level evidence), etc. Providing such explanations helps users understand, verify, and trust the detector’s decisions.
Cross-Model Generalizability. Given the rapid evolution of text-to-image generative models, new architectures and generated images are continually emerging, which may degrade the performance of detectors trained on previous generative models. Therefore, the detector should be robust to various model iterations and new architectures, and capable of leveraging generalizable evidence to assess image provenance, rather than relying solely on model-specific features.
Training-Free Design. Furthermore, traditional AI image detection frameworks typically depend on collecting large-scale training datasets of AI-generated images. As generative models evolve, these frameworks require continual updates to their training data, incurring significant maintenance costs. In contrast, the core of our detector is designed to be training-free: it does not require any dedicated training set, and can be directly applied to detect AI-generated images from any model, including those unseen during development.
Detector’s Capabilities
Our detector is assumed to have access to the following resources and conditions:
  • Image-File Access. The detector can access the image file itself, including its metadata if available (e.g., camera details, timestamps, and editing history).
  • External Forensic Tools. The detector is permitted to utilize a suite of external forensic tools to analyze the image and gather supporting evidence. This includes pretrained AI detection models, reverse image search engines, and vision-language models.
The AIFo Framework
Design Rationale
Unlike existing approaches that rely solely on pre-trained classifiers or VLMs, our main design principle is to emulate the procedural reasoning processes employed by human forensic experts [22, 38], while still leveraging the strengths of existing approaches. To this end, we leverage LLMs as autonomous agents to develop an agentic framework that incorporates the key capabilities of human forensic experts. First, the framework is designed to employ and coordinate a suite of forensic tools, such as pre-trained classifiers, metadata extraction, reverse image search, and VLM-based analysis. Second, rather than relying on a single fixed source of evidence, the framework should be able to cross-check evidence from multiple tools before forming the final decision. Next, when the collected evidence indicates conflicting signals, the framework is expected to resolve these conflicts through human-like reasoning activities, such as assessing the reliability of evidence sources and engaging in structured debate among the expert team. The core framework is training-free, building on top of pre-trained classifiers and state-of-the-art VLMs as forensic tools. At the same time, an optional memory module allows the framework to accumulate historical cases and progressively improve performance. Altogether, we aim to build a training-free, interpretable, and robust framework for AI-generated image detection, which not only leverages existing detection approaches but also possesses key capabilities of human forensic experts, e.g., integrating various evidence, evaluating their reliability, and resolving conflicts among them.
Overview
Figure 1: High-level overview of our proposed AIFo.
Our proposed framework, Agent-based Image Forensics (AIFo), is a training-free, LLM-based multi-agent system for AI-generated image detection, designed to mimic the procedural reasoning workflow of human forensic experts. As illustrated in Figure 1, the core of our framework [^1] We introduce the optional memory module in Section 5. consists of a Toolbox $\mathcal{T}=\{T_{1},T_{2},\dots,T_{n}\}$ containing $n$ different forensic tools, an Evidence Gatherer Agent $A_{EG}$, a Reasoning Agent $A_{R}$, two Debate Agent $A_{D1}$ and $A_{D2}$ and a Judge Agent $A_{J}$. Given an input image of unknown authenticity from the set of all images, $I\in\mathcal{V}$, the Evidence Gatherer $A_{1}$ first employs tools from the Toolbox. Specifically, the Toolbox includes four main types of forensic tools: (1) Image Reverse Search to locate visually similar images and retrieve their provenance information from online sources; (2) Metadata Extraction to recover embedded EXIF or generated metadata from the image file; (3) VLM-Based Analysis to perform semantic-level reasoning and contextual understanding of the image leveraging VLMs; and (4) Pre-Trained Classifiers to apply pre-trained models for binary classification of the image.
Then, the Evidence Gatherer Agent executes each tool $T_{i}\in\mathcal{T}$ to produce a piece of evidence:
$e_{i}=T_{i}(I),\quad T_{i}:\mathcal{V}\rightarrow\mathcal{E}_{i},$
where $\mathcal{E}_{i}$ denotes the evidence space for tool $T_{i}$. For example, the metadata extraction tool may return detailed EXIF information embedded in the image file, such as "EXIF:Model": "Canon EOS 5D Mark IV" or "EXIF:LensModel": "EF24-70mm f/2.8L II USM", which can serve as important evidence for authenticity assessment. The complete evidence set is:
$E=\{e_{i}\mid T_{i}\in\mathcal{T}\}.$
Next, the Reasoning Agent $A_{R}$ examines the entire evidence set $\mathcal{E}$ and determines whether the evidence is sufficient and consistent enough to make a reliable decision. Evidence sufficiency refers to whether the collected evidence covers enough aspects of the input image, e.g., without missing key metadata or reverse search results. Evidence consistency measures the extent to which different pieces of evidence point toward the same conclusion, e.g., when the majority of evidence supports that the image is AI-generated. If the evidence is considered sufficient and consistent enough, the Reasoning Agent generates a final judgment $D\in\mathcal{D}$ and a human-readable explanation $R\in\mathcal{R}$:
$A_{R}:\mathcal{E}\xrightarrow{}\mathcal{D}\times\mathcal{R}.$
Otherwise, the system initiates a multi-round debate process. Two Debate Agents $A_{D1}$ and $A_{D2}$ take opposing stances (pro vs. contra) regarding whether the image is AI-generated or real. They exchange arguments over $n$ rounds, and a Judge Agent $A_{J}$ observes the debate history $H$ together with the tool-derived evidence $\mathcal{E}$, and produces the final judgment $D\in\mathcal{D}$ along with an explanation $R\in\mathcal{R}$:
$A_{J}:(H,\mathcal{E})\rightarrow\mathcal{D}\times\mathcal{R}.$
This hierarchical reasoning-debate framework ensures that decisions are made either directly from sufficient evidence or through a careful debating process that simulates human forensic reasoning when the evidence is inconclusive or conflicting.
LLM-Based Agents
The LLM-based agents act as the central reasoning and orchestration units of our multi-agent framework, responsible for applying forensic tools, evaluating the reliability of collected evidence, and making decisions based on existing evidence. All agents are instantiated from a LLM (GPT-4.1-2025-04-14), with specialized prompts defining their respective roles and responsibilities. The detailed prompts used for each agent are provided in Appendix A.
The Evidence Gatherer has full access to the Toolbox and is responsible for invoking the forensic tools to collect evidence. Its decision-making process is guided by a carefully engineered prompt that contains three critical components: (1) a clear definition of the agent’s task, which is acting as an AI image forensics expert to determine whether the input image is AI-generated or real; (2) rigorous definitions of “AI-generated image” and “real image” (see Section 2.1) to ensure consistent interpretation during analysis; (3) a comprehensive description of all available tools, including their capabilities. With these explicit task definitions and resource descriptions, the LLM gains both the contextual understanding and the operational awareness to invoke the tools and collect results.
The Reasoning Agent is responsible for conducting an initial assessment of the evidence collected by the evidence gatherer. It first evaluates whether the evidence set is sufficient and consistent enough. This evidence set check is implemented via a dedicated prompt, which instructs the LLM to internally assess the completeness (sufficiency) and consistency of the evidence and return a binary decision (true / false). If true, the reasoning agent proceeds to generate a final judgment and an explanation with another prompt that: (1) defines its responsibility to synthesize the collected evidence and produce a final binary judgment (AI-generated or real) with an accompanying explanation; (2) instructs the agent to evaluate the reliability of each evidence source, taking into account the confidence level, potential biases, and the trustworthiness of the generating tool; (3) emphasizes the importance of generating a human-readable, logically structured reasoning chain. If the evidence is insufficient or conflicting, the reasoning agent abstains from producing a final decision and instead triggers the multi-agent debate mechanism.
The multi-agent debate mechanism is designed to resolve ambiguity and enhance the robustness of the final judgment when the evidence is inconclusive or conflicting.
The Debate Agents are responsible for engaging in a structured multi-round debate when the evidence is deemed insufficient or inconsistent. The debate process lasts up to $n$ rounds. In the first round, the two debate agents are guided by designed prompts to take opposing stances: one agent provides arguments supporting that the image is AI-generated based on collected evidence, while the other provides arguments supporting that the image is real. In subsequent rounds, both agents are instructed to refine or strengthen their reasoning based on the arguments made by the opposing side in the previous round. Through this iterative exchange, the debate agents progressively sharpen their analyses and solve potential conflicts in the evidence interpretation.
A Judge Agent is responsible for supervising the debate process. At the end of each round, the Judge Agent internally evaluates whether the debate has reached sufficient coverage and clarity to make a final decision. It may decide to terminate the debate early if the arguments are deemed sufficient, or allow the debate to proceed for additional rounds if further clarification is needed. Once the debate concludes, the Judge Agent generates the final judgment and an explanation. Unlike the Reasoning Agent, which relies solely on the collected evidence, the Judge Agent is guided to synthesize both the tool-derived evidence and the debate history. This design ensures that the final decision integrates multiple sources of evidence with structured deliberation, leading to a transparent and defensible forensic conclusion.
Forensic Tools
The forensic Toolbox is a collection of specialized modules that the Evidence Gatherer can invoke to analyze the input image. The tools can be broadly categorized into the following four classes:
  • Reverse Image Search Tools: Assess image provenance by querying external online sources.
  • Metadata Extraction Tool: Parse metadata information from the image file.
  • Pre-Trained Classifiers Tool: Apply static models to assess authenticity.
  • VLM-Based Reasoning Tool: Perform semantic-level analysis using VLMs.
First, reverse image search tools are designed to identify the provenance and distribution history of input images by searching for exact matches or visually similar content across the Internet. These tools are particularly useful for determining image provenance, e.g., AI-generated images may appear on generative art platforms, while real images are more likely to be found on reputable news or photographic sites. We implement two complementary reverse image search tools. The first tool leverages the Google Cloud Vision API’s Web Detection service [7], which provides programmatic access to Google’s extensive image indexing capabilities. It primarily returns exact matches of the query image by analyzing its visual features and comparing them against Google’s web-scale image database. The API outputs structured data such as webpage titles, URLs, and contextual information where identical or highly similar images are found. However, the returned information can sometimes be sparse or insufficient for reliable judgment. To address this limitation and broaden the search scope, we implement a second approach inspired by Xu et al. [55] that simulates authentic human user behavior through web interface automation. This tool directly interacts with the Google Images search interface using Playwright, [^2] https://playwright.dev/ a modern web automation framework, to perform searches beyond the API’s restricted results. Unlike the first tool, this approach captures not only exact matches but also a wider range of visually similar images, thereby providing richer contextual evidence. It automatically uploads the input image, captures the results section from the returned webpage while filtering out advertisements and noise, and summarizes key information such as image match identification and the provenance of similar images.
Metadata analysis tool focuses on extracting and analyzing technical metadata embedded within image files to identify authenticity markers that distinguish real photographs from AI-generated content. Digital cameras and imaging devices typically embed rich metadata (EXIF data) including camera settings, GPS coordinates, timestamps, and device-specific information that are often absent or inconsistent in synthetically generated images. For example, as shown in Table 1, a genuine photograph may contain entries like EXIF:Make and EXIF:Model indicating the camera manufacturer and model. They may also include realistic optical parameters or position data such as EXIF:LensInfo and Composite:GPSPosition that are indicative of a real-world capture event. In contrast, AI-generated images frequently lack such detailed metadata or contain synthetic traces. For instance, some AI images explicitly embed generation information or even the original text prompt used for synthesis, e.g., EXIF UserComment. These anomalies serve as strong indicators of AI generation.
Our metadata extraction tool employs ExifTool, [^3] https://exiftool.org/ a comprehensive metadata manipulation library, to perform deep analysis of image files. Rather than extracting all available metadata fields, which can be overwhelming and include irrelevant information, our system implements a selective extraction strategy based on two key filtering mechanisms. First, we maintain a curated list of exact-match key fields $\mathrm{KEY\_FIELD\_EXACT}$ that have been empirically determined to provide strong authenticity signals. These include critical camera parameters such as focal length, aperture settings, ISO values, and camera manufacturer information. Second, we implement prefix-based filtering $\mathrm{KEY\_FIELD\_PREFIXES}$ to capture related metadata families. For example, all fields for camera manufacturer custom information are captured through the “MakerNotes” prefix. This dual-filtering approach ensures comprehensive coverage of relevant metadata while filtering out noise. The complete list of all key fields and prefixes is provided in the Appendix B.
Table 1: Comparison of metadata signals for distinguishing real and AI-generated images.
Metadata Field Example Value
Real Image Signal
EXIF:Make Canon
EXIF:LensInfo 4.1 123 3.5 6.4
Composite:GPSPosition 28.35 N, 81.59 W
AI Image Signal
JUMBF:Description AI Generated Image
EXIF:UserComment "A full body of a cat..."
Pre-trained classifiers tool leverages pre-trained models specifically fine-tuned for AI-generated image classification. These models have been trained on diverse datasets containing both authentic and AI-generated images, enabling them to learn discriminative features that distinguish between the two categories. We select the top five most downloaded classification models for AI-generated image detection available on Hugging Face [1]:
  • haywoodsloan/ai-image-detector-deploy [3]
  • Organika/sdxl-detector [12]
  • legekka/AI-Anime-Image-Detector-ViT [4]
  • Smogy/SMOGY-Ai-images-detector [13]
  • NYUAD-ComNets/NYUAD_AI-generated_images
    _detector
     [11]
Our implementation employs an ensemble approach that integrates multiple transformer models to enhance detection robustness and accuracy. The ensemble includes models trained on different synthetic image generation techniques, ensuring comprehensive coverage of various AI generation paradigms including GANs, diffusion models, and other neural synthesis methods. Each transformer model performs independent classification, producing probability distributions over the binary classification space. The final prediction score is obtained through a weighted voting mechanism, where each model’s contribution is scaled by its weight parameter. This weighted ensemble score is computed as:
$\text{Prediction Score}=\frac{\sum_{i=1}^{N}w_{i}\cdot s_{i}}{\sum_{i=1}^{N}w_{i}},$
where $w_{i}$ represents the weight of the $i$-th model $\theta_{i}$, $s_{i}$ is the AI confidence score from $\theta_{i}$, and $N$ is the total number of loaded models. Each model is assigned an equal weight $w_{i}=1$ to ensure balanced contribution and avoid over-reliance on any individual model during inference.
VLM-based reasoning tools employs VLMs to perform sophisticated visual analysis. These models leverage their extensive training on diverse image-text pairs to identify subtle visual patterns, artifacts, and inconsistencies that may indicate AI generation. Our implementation utilizes GPT-4.1 [8] to conduct detailed image analysis. A crafted prompt guides the model to focus on specific visual characteristics that are indicative of AI generation versus authentic photography. The complete prompt is provided in Table 15 in the Appendix. When the model identifies an image as AI-generated, it is prompted to provide detailed evidence in the form of specific visual artifacts. These include unnatural textures or patterns that deviate from expected material properties, inconsistent lighting or shadow directions that violate physical principles, anatomical errors in human or animal subjects, unusual distortions in object boundaries, text rendering abnormalities, symmetry issues, and contextual inconsistencies in background elements. For images classified as real, the model is instructed to explain the supporting visual characteristics, such as realistic anatomical proportions and coherent environmental context. Unlike the other forensic tools, VLMs not only enable binary classification but also provide detailed justifications and fine-grained visual analyses that enhance the interpretability of the final decision. Additionally, the model provides a confidence assessment (high, medium, or low) based on the strength and clarity of the observed evidence. The deterministic configuration (temperature = 0, seed = 42) ensures consistent and reproducible analysis results across multiple runs.
Evaluation
Dataset Construction
To comprehensively evaluate our agentic framework, we consider both controlled and real-world scenarios. We therefore construct our benchmark dataset to cover two distinct settings: in-the-lab and in-the-wild. The in-the-lab setting consists of images collected from well-curated, controlled datasets commonly used in prior research, while the in-the-wild setting comprises images sourced from diverse, unconstrained online platforms, reflecting the complexity and unpredictability of real-world data.
Our curated dataset comprises a total of 6000 images, evenly distributed across the two settings, with each containing 1500 AI-generated images and 1500 real images. For the in-the-lab setting, real images are sampled from well-established, curated datasets frequently used in image generation and detection research. Specifically, we sample 500 images each from Flickr30k [40], ImageNet [20], and the DIV2K dataset [16]. The AI-generated counterparts are obtained from GenImage [61] and FakeBench [33]. We sample 100 images from each of the 8 generative models included in GenImage, and 70 images from each of the 10 models in FakeBench, resulting in a total of 1500 AI-generated images.
[…]
Table 2: Dataset construction statistics.
Source Datasets / Platforms Type # Images
In-the-lab Setting
Flickr30k [40] Real Image 500
ImageNet [20] Real Image 500
DIV2K [16] Real Image 500
GenImage [61] AI Image 800
FakeBench [33] AI Image 700
In-the-wild Setting
Holopix50k [26] Real Image 500
Flickr [6] Real Image 500
Wikimedia Commons [14] Real Image 500
Lexica [9] AI Image 500
NightCafe [10] AI Image 500
Civitai [5] AI Image 500
Experimental Setup
Implementation Details. To evaluate the performance of our AIFo framework, we conduct experiments on the constructed benchmark dataset, comparing our agentic approach against a range of baseline methods. Our framework is implemented using LangGraph, [^4] https://langchain-ai.github.io/langgraph/ which is specifically designed for building stateful, multi-agent applications with LLMs. We select GPT-4.1 (gpt-4.1-2025-04-14 version) as the backbone for our LLM agents, and we set the temperature to 0 and the seed to 42 to ensure deterministic behavior across runs.
Evaluation Metrics. We formulate the evaluation as a standard binary classification task, where the goal is to assess the framework’s ability to distinguish between AI-generated and real images. Specifically, we define correctly identifying an AI-generated image as a True Positive (TP) and correctly identifying a real image as a True Negative (TN). Based on these definitions, we compute the following standard evaluation metrics: Accuracy, Precision, Recall, and F1-score. In addition to evaluating the framework’s overall performance on the entire dataset, we further analyze its behavior separately on the in-the-lab and in-the-wild subsets to assess its robustness across controlled and real-world scenarios.
Baseline Methods. To comprehensively assess the effectiveness of our proposed multi-agent framework, we compare it against a diverse set of representative baseline methods spanning both conventional and vision-language modeling paradigms. For traditional classifier-based approaches, we include CNNSpot [53], DE-FAKE [46], and PatchCraft [59], with their details discussed in Section 7.1. For VLM baselines, we adopt state-of-the-art multimodal models including GPT-4o and GPT-4.1, which have demonstrated advanced visual reasoning capabilities. During evaluation, each model is provided with the input image along with the prompt Is this a fake or real image?, and instructed to return a structured binary classification result.
Results
Table 3: Performance comparison of different methods on our benchmark dataset comprising three evaluation subsets: Overall, In-the-Lab, and In-the-Wild. Metrics reported are Accuracy (Acc), Precision (Prec), Recall (Rec), and F1-score (F1). Best results are highlighted in bold and second best results are underlined.
Overall In-the-Lab In-the-Wild
Method Acc Prec Rec F1 Acc Prec Rec F1 Acc Prec Rec F1
CNNSpot [53] 0.5277 0.9826 0.0563 0.1066 0.5553 0.9882 0.1120 0.2012 0.5000 0.5000 0.0007 0.0013
PatchCraft [59] 0.6517 0.7423 0.4647 0.5715 0.8123 0.8704 0.7340 0.7964 0.4910 0.4780 0.1953 0.2773
DE-FAKE [46] 0.7142 0.6820 0.8027 0.7374 0.6720 0.6673 0.6860 0.6765 0.7563 0.6933 0.9193 0.7905
GPT-4.1 [8] 0.9416 0.9913 0.8910 0.9385 0.9332 0.9932 0.8723 0.9288 0.9500 0.9895 0.9097 0.9479
GPT-4o [39] 0.9483 0.9920 0.9038 0.9458 0.9537 0.9938 0.9130 0.9517 0.9428 0.9900 0.8947 0.9399
AIFo w/o debate (ours) 0.9635 0.9922 0.9343 0.9624 0.9730 0.9924 0.9533 0.9725 0.9540 0.9921 0.9153 0.9521
AIFo (ours) 0.9705 0.9920 0.9487 0.9698 0.9740 0.9917 0.9560 0.9735 0.9670 0.9923 0.9413 0.9661
Table 3 presents the comprehensive performance comparison of our multi-agent framework against five baseline methods across three evaluation settings. A detailed breakdown of their performance is provided in Table 17 in the Appendix. These results demonstrate the superior effectiveness of our proposed approach in AI-generated image detection under both controlled and real-world conditions.
Traditional Classifier Baselines. The traditional methods show varied performance patterns. CNNSpot [53] exhibits poor overall detection performance, close to random guessing. Its extremely low recall and F1-score indicate that it fails to effectively detect AI-generated images. DE-FAKE [46] achieves moderate performance, with an accuracy of 0.7142 and an F1 score of 0.7374, slightly outperforming PatchCraft [59]. These results highlight the inherent limitations of traditional classification methods in coping with rapidly evolving and continuously updated diverse AI-generated content in complex real-world scenarios.
Vision-Language Model Baselines. Both GPT-4.1 and GPT-4o demonstrate strong performance, with accuracy exceeding 0.94 and F1-scores above 0.93. However, the recall rates for both models are relatively lower compared to their precision, indicating that they still miss a significant portion of AI-generated images. This suggests challenges remain for these models in detecting certain types of AI-generated content. Our agent framework addresses this limitation through the multi-agent architecture, which combines the strengths of multiple specialized tools to achieve more reliable detection.
AIFo Framework Performance. Compared to all baseline methods, our AIFo framework consistently achieves the best performance across almost all metrics and evaluation settings. In the overall evaluation, AIFo attains an accuracy of 0.9705 and an F1-score of 0.9698, surpassing the strongest baseline GPT-4o by absolute margins of 2.22% in accuracy and 2.40% in F1-score. Moreover, this improvement is observed consistently in both in-the-lab and in-the-wild data, which span over 20 generative models, demonstrating the strong generalization capability of our framework. Through a training-free paradigm, our method is minimally affected by the evolution and iteration of generative models.
Impact of Multi-Agent Debate Mechanism. We also evaluate the impact of disabling the multi-agent debate mechanism on the performance. Notably, even without the debate mechanism, our AIFo framework still significantly outperforms GPT-4o. With the debate mechanism incorporated, our framework’s accuracy is further improved, particularly in in-the-wild setting, where it surpasses GPT-4o by 2.4%. These results demonstrate that the debate mechanism serves as an effective refinement stage, resolving potential reasoning uncertainties caused by conflicting evidence and enhancing the reliability of the final decision.
Laboratory vs. Wild Environment Analysis. A comparison between controlled (in-the-lab) and real-world (in-the-wild) environments reveals key differences. Traditional classifiers often fail to generalize: for example, PatchCraft’s F1-score drops from 0.7964 in the laboratory to 0.2773 in the wild, and CNNSpot nearly collapses entirely. This highlights the difficulty of transferring models trained on curated datasets to diverse real-world content. In contrast, our agent framework achieves consistently strong results, with F1-scores of 0.9735 in the laboratory and 0.9661 in the wild, across more than 20 generative models. These findings confirm the framework’s ability to generalize beyond controlled datasets and effectively handle the diversity of real-world AI-generated content.
Quantitative and Qualitative Analysis. To better understand where AIFo achieves improvements over GPT-4o, we analyze all samples misclassified by GPT-4o but correctly identified by AIFo. We find that AIFo correctly identifies 136 more AI-generated images compared to GPT-4o out of the total 6,000 samples. Among these corrected cases, approximately 38% are attributed to decisive reverse image search evidence, primarily from in-the-wild sources. Another 35% are resolved through metadata signals, which prove particularly effective for images containing detailed camera parameters or generation prompts embedded by community diffusion models. The remaining 27% rely on transformer-based classifier results as supporting evidence. In approximately 30% of all corrected cases, AIFo invokes the debate mechanism to resolve ambiguous or conflicting evidence, ultimately overturning the initial VLM-based assessment.
To further demonstrate the effectiveness of AIFo, we present qualitative examples in Figure 2 that show how the system integrates multiple sources of evidence and, when necessary, employs a debate mechanism to reach accurate conclusions. In the first real-image case, the Evidence Gatherer agent successfully collects evidence from four forensic tools, including image sources, camera parameters, VLM analysis, and classifier scores. The Reasoning Agent judges the evidence sufficiency and, by synthesizing all sources, correctly concludes that the image is real.
In the second AI-generated case, the realistic appearance misleads the baseline GPT-4o and the VLM tool, but other tools indicated an AI origin, creating conflicting signals. In this case, the Reasoning Agent judges that the evidence is inconsistent for a direct decision and instead triggers the debate mechanism. The pro-debate agent and con-debate agent present arguments supporting their respective stances, and the decisive argument came from the pro side, which highlighted the exact match of the image found on an AI platform. Ultimately, the Judge Agent weighs the conflicting evidence and reaches the correct conclusion that the image is AI-generated.
These cases highlight two key strengths of our framework: its ability to override misleading visual evidence using stronger provenance signals, and its capacity to reconcile conflicting evidence through structured debate. These capabilities are essential for reliable detection in both benchmarks and real-world scenarios.
Figure 2: Examples of our agent framework’s decision-making process, demonstrating diverse evidence integration across different image types and sources.
Tool Reliability and Decision Pattern Analysis. To gain deeper insights into the internal decision-making processes of our multi-agent framework, we conduct a quantitative analysis of the reliability and usage patterns of individual tools, as illustrated in Figure 3. We define a consistent judgment as a case in which the agent’s final decision is in agreement with the evidence provided by a given tool. The reliability rate of a tool is then computed as the ratio of its consistent judgments to the total number of valid (i.e., informative) pieces of evidence it produces. A higher reliability rate indicates that the agent frequently trusts and aligns with that tool’s output, suggesting its potential influence on the overall decision-making process. We also compute the coverage rate of each tool, defined as the proportion of cases in which the tool provides valid evidence relative to the total number of agent decisions.
From 3(a) and 3(b), metadata extraction and the first reverse search tool focused on exact matches achieve the highest reliability rates, meaning their evidence plays a significant role in shaping the agent’s final decisions when available. However, their lower coverage rates show they are only relevant to a subset of images. This is because in our benchmark dataset, only a small fraction of images contain rich EXIF metadata or have provenance information retrievable using the first reverse search tool. These tools thus act as high-precision, low-frequency decision anchors, exerting strong influence when available but not contributing in all cases. In contrast, the VLM analysis tool, pre-trained classifiers, and the second reverse search tool specialized in similar image retrieval achieve nearly full coverage. While the VLM’s visual reasoning remains relatively reliable, the outputs of the second reverse search tool and the classifiers are less dependable. This suggests that the agent will internally assign lower weights to these pieces of evidence. Hence, these tools serve as broad-coverage, moderate-confidence sources of evidence, providing additional context and support for the final decision but not dominating the reasoning process.
Overall, the agent appears to follow a tiered weighting strategy: prioritizing high-reliability tools when available, while leveraging high-coverage tools to maintain decision robustness.
(a) Tool reliability rates
(b) Tool coverage rates
Figure 3: Analysis of individual tool contributions to the agent framework: (a) reliability rates measuring agent trust in each tool’s evidence, and (b) coverage rates showing the proportion of decisions where each tool provides informative evidence.
Inference Efficiency and Cost Analysis. While AIFo achieves high detection accuracy, its architecture introduces additional computational cost compared to single model such as GPT-4o. To assess the efficiency of our framework, we measure the end-to-end latency and token usage for processing a single image. Table 4 summarizes the inference latency and token consumption per image for different methods. On average, AIFo requires 40.08 seconds per image, approximately 7.5 times slower than GPT-4o’s 5.31 seconds. The token usage also increases significantly, with AIFo consuming an average of 5230.86 tokens per image, compared to GPT-4o’s 715.05 tokens. The majority of the additional latency and token usage originates from the multi-round llm invocation and the execution of external tools. Despite the increased computational cost, AIFo achieves a consistent 2-3 % improvement in accuracy and over 4 % gain in recall. Meanwhile, it is important to note that AIFo operates in a training-free manner, eliminating the need for costly model retraining or fine-tuning as generative models evolve. Most importantly, unlike single model reasoning that relies solely on visual analysis, our framework integrates diverse sources of evidence, offering more verifiable interpretability that aligns with human forensic reasoning.
Table 4: Average inference latency and token usage per image.
Method Avg. Latency Avg. Token Usage
GPT-4o 5.31s 715.05
AIFo (w/o debate) 25.43s 2728.29
AIFo 40.08s 5230.86
Robustness Analysis
To assess the robustness of our multi-agent detection framework under realistic conditions, we evaluate its performance on perturbed versions of the dataset. The goal of this analysis is to determine whether the framework can maintain consistent accuracy when the input images undergo quality degradation. We apply a series of standard image degradation techniques, including: Blurring. We apply Gaussian blur with a fixed radius of 2, simulating defocus or motion blur; Sharpening. We enhance image sharpness using a sharpening factor of 2.0, which may alter local edge distributions; Gaussian Noise. We add Gaussian noise with mean 0 and variance 2 to simulate sensor-level or environmental noise. These image transformations represent common distortions encountered in real-world image acquisition or compression pipelines. For each perturbation type, we measure the detection performance on the perturbed images and compare the results against the best baseline, GPT-4o.
Table 5 presents the comprehensive performance comparison between our AIFo framework and the GPT-4o baseline across three types of image perturbations: Blurring (Blu), Gaussian noise (Noi), and sharpening (Sha). The results demonstrate that our framework maintains superior performance under these robustness testing conditions.
Table 5: Performance comparison of AIFo vs GPT-4o under different image perturbations on the overall dataset. AIFo superior results are highlighted in bold.
GPT-4o AIFo (Ours)
Acc Prec Rec F1 Acc Prec Rec F1
Blu 0.8818 0.9662 0.7913 0.8701 0.9047 0.9380 0.8667 0.9009
Noi 0.9462 0.9866 0.9047 0.9438 0.9690 0.9879 0.9497 0.9684
Sha 0.9410 0.9926 0.8887 0.9377 0.9670 0.9902 0.9433 0.9662
Our AIFo framework consistently outperforms GPT-4o across all perturbation types and most evaluation metrics. Under blurring conditions, both GPT-4o and AIFo experience noticeable performance degradation, as visual analysis tools are inherently sensitive to loss of a great amount of image pixel features. Nevertheless, AIFo still achieves an accuracy of 0.9047, outperforming GPT-4o’s 0.8818. In contrast, under noisy and sharpened conditions, AIFo demonstrates remarkable robustness: its accuracy remains at 0.9710 and 0.9670 respectively, with only minimal decline compared to the clean setting, and consistently surpasses GPT-4o. This robustness can be attributed to the fact that, apart from pre-trained classifiers, the majority of tools integrated in our framework are largely insensitive to the perturbations, thereby stabilizing overall performance.
Evasive Attack Analysis
To evaluate the resilience of AIFo against evasive attacks, we simulate two representative attack scenarios that target the framework’s key evidence sources. The first attack employs reverse image search manipulation techniques and the second attack involves metadata forgery. In the first scenario, we manipulate the provenance evidence returned by the reverse image search tool. Specifically, we use GPT-4o to generate counterfactual search results for each image. For real images, we fabricate search results indicating that the image was sourced from an AI generation platform, while for AI-generated images, we create results suggesting the image originated from a reputable photography website. These results are then injected into the evidence set returned to the agent. In the second scenario, we perform metadata forgery by swapping EXIF metadata between AI-generated and real images. Real images are randomly assigned metadata extracted from AI-generated samples, while AI-generated images receive real images’ metadata.
As shown in Table 6, AIFo experiences a moderate performance degradation under both evasive attack settings. This degradation is primarily due to the framework’s limitation in verifying the authenticity of evidence returned by external tools. Since AIFo is designed to treat tool outputs as trustworthy forensic sources, deliberately falsified information can mislead the agent, resulting in false judgments. To enhance robustness against such attacks, several potential defenses can be considered. Firstly, implementing cross-tool consistency validation can help identify conflicting evidence that may indicate manipulation such as verifying that metadata timestamps and reverse-search provenance sources are mutually coherent. Secondly, trust-weighted evidence aggregation can be introduced, where each tool’s output is dynamically weighted based on historical reliability. Finally, incorporating external verification layers such as digital watermark authentication can help validate evidence before feeding it into the reasoning pipeline. These strategies would allow AIFo to better distinguish adversarially manipulated evidence, thereby improving its resilience against real-world evasive attacks.
Table 6: Performance of AIFo under two evasive attack scenarios.
Attack Type Acc Prec Rec F1
None (clean) 0.9705 0.9920 0.9487 0.9698
Reverse-search manipulation 0.8971 0.8690 0.9353 0.9010
Metadata forgery 0.8702 0.8399 0.9147 0.8757
Ablation Study
To investigate the individual contributions of different tool categories within our multi-agent framework, we conduct ablation experiments to quantify the impact of each of the four tool types on overall detection performance. To assess the importance of each category, we adopt a leave-one-out strategy: in each ablation run, we disable one tool category while keeping the others unchanged, and evaluate the resulting performance on the full benchmark dataset. Meanwhile, we temporarily disable the debate mechanism to better isolate and focus on the contributions of the tools themselves. By observing the performance drop associated with the removal of each tool category, we can quantify its relative contribution to the framework’s overall effectiveness. This analysis provides insight into which components are most essential for robust and reliable detection, and helps guide future efforts in optimizing or simplifying the system without substantial loss in performance.
Figure 4 presents the performance degradation observed when each tool category is disabled. The results clearly demonstrate that the vision-language model (VLM) tool contributes the most critical evidence to the overall detection pipeline. Removing the VLM tool causes the largest performance degradation across all metrics, with accuracy dropping below 0.85, and recall falling below 0.70. This indicates that direct image analysis from state-of-the-art VLMs remains the main source of evidence for the reasoning agent.
Disabling any of the metadata extraction, reverse image search, or pre-trained classifier modules leads to a slight decrease in both accuracy and F1 score. This indicates that the evidence provided by these tools offers complementary information that enhances the capability of the reasoning agent. Notably, even without these modules, our framework still surpasses the GPT-4o baseline, confirming that these tools contribute complementary signals that strengthen the robustness of the overall reasoning process.
Overall, the ablation study validates the necessity of a diverse Toolbox for reliable detection. These findings highlight that performance gains in our multi-agent framework arise not from a single dominant component, but from the effective integration of cross-source and complementary evidence sources.
Figure 4: Performance degradation when each tool is disabled from the framework.
Takeaway
Our evaluation demonstrates the advantages of the proposed multi-agent framework: it consistently achieves the highest detection performance on five benchmark datasets in the lab and in-the-wild images from six online platforms, covering AI-generated content produced by at least 20 models. The framework also shows stronger robustness than the best baseline, GPT-4o, against three common perturbations. We find that our framework can dynamically weigh the confidence of evidence gathered from each tool: when metadata and reverse search tools provide valid evidence, it tends to prioritize them; otherwise, it falls back on VLM analysis and traditional classifiers. The debate module further addresses conflicting evidence and prevents the framework from over-relying on a single source of evidence. Altogether, the framework demonstrates greater effectiveness and robustness compared to single-model approaches for AI-generated image detection.
Memory-Augmented Reasoning
Motivation and Design
Figure 5: Overview of the memory-augmented reasoning module.
Our current designed detectors operate as stateless processes, analyzing each image independently without leveraging historical context or learning from past decisions. However, forensic analysis is inherently iterative and cumulative, where insights from previous cases can inform future judgments [25]. To address this limitation, we introduce a memory-augmented reasoning module that enables the framework to maintain a knowledge base of past cases and decisions. As illustrated in Figure 5, the module captures, indexes, and retrieves relevant historical context to inform current detection decisions. The designed knowledge base systematically collects both successful and failed detection cases, storing image embedding vectors together with their associated classification results, collected evidence, and reasoning processes. For failed cases, it additionally stores a reflective analysis that captures the causes of misclassification. When analyzing a new image, the system performs similarity-based retrieval over this knowledge base to identify relevant historical cases.
We hypothesize that the introduction of similar historical cases enables the agent to learn from experience and potentially adjust the weights of evidence from different forensic tools based on past performance patterns. Therefore, the enhanced system first queries its accumulated experience to identify potentially relevant historical cases rather than approaching each image as an isolated classification problem.
Implementation Details
Knowledge Base Dataset. To avoid data leakage, we construct the knowledge base image repository using a completely separate set of images from those used in the main benchmark dataset described in  Section 4.1. Following the same collection and sampling methodology as the benchmark dataset, we curate 600 new images for the knowledge base, containing 300 AI images and 300 real images.
Building Knowledge Base. Our knowledge base system is built on a hybrid architecture that combines vector similarity search for content retrieval with structured metadata storage for case management and analysis. The implementation leverages CLIP [42] embeddings to capture high-level semantic and visual features that enable effective similarity matching across images. Each image processed by the framework undergoes feature extraction using the CLIP-ViT-B/32 model, generating 512-dimensional dense vector representations. These embeddings serve as the primary indexing mechanism, enabling efficient similarity-based retrieval of historically relevant cases. The knowledge base maintains separate indices for successful and failed detection cases. For successful cases, the system records the complete set of evidence used in decision-making and the full analytical process. For failed cases, the system records all the above information and additionally stores a structured reflection, providing deeper insight into the causes of misclassification. The reflection is generated by GPT-4.1 to examine the complete analytical pathway and identify potential failure points. This process produces insights covering evidence misinterpretation, tool reliability assessment, and reasoning inconsistencies. These generated reflections create a systematic learning mechanism that helps the system avoid repeating similar analytical errors. Finally, the knowledge base module is integrated seamlessly as one extra forensic tool. During the evidence collection phase, the module is activated.
Inference. Upon receiving a new image, the system computes its CLIP embedding and performs a similarity search over the knowledge base indices to retrieve the top-$k$ most relevant historical cases, with defaults of $k=1$ and a similarity threshold of $0.85$ to ensure only sufficiently close matches are returned. The retrieved cases are then passed to the reasoning agent as additional context, allowing it to leverage past experiences to inform the current analysis.
Results Analysis
We did not incorporate the memory module in the main evaluation, as it requires ground-truth labels during the experience accumulation phase. Allowing such label-dependent learning at test time would introduce unfair advantages to our framework, compared to baselines that do not access these labels. Instead, to highlight the potential of the memory module, we directly focus on failure cases from the main evaluation. These are the scenarios where the knowledge base and additional experience are most needed.
To more directly assess its effectiveness, we conduct a targeted evaluation focusing on failure cases from the original framework. Specifically, we identify 50 misclassified images from the benchmark set for which the knowledge base contained semantically similar counterparts. We then re-evaluated these images with the knowledge base module enabled, examining whether the additional contextual evidence could help correct the prior misjudgments. Table 7 summarizes the results by different image types and settings, showing the number of errors successfully corrected after incorporating knowledge base evidence. The results demonstrate that over 40% of error cases are successfully corrected when the memory module is employed. Furthermore, the example illustrated in Figure 6 highlights how retrieved similar cases influenced the reasoning process. When presented with similar past cases, the agents internally reweighted the credibility of different forensic tools, reinterpreted conflicting evidence, and ultimately arrived at the correct conclusion. These observations suggest that the memory-augmented reasoning offers substantial benefits in failure recovery and demonstrates strong potential for enhancing adaptive reasoning in forensic scenarios.
Table 7: Number of errors before and after incorporating similar case history. FP: real images misclassified as AI images; FN: AI images misclassified as real images.
In-the-Lab In-the-Wild
FP FN FP FN Total
Before 2 12 1 35 50
After 1 6 0 22 29
Figure 6: Example of memory-augmented reasoning showing how historical cases influence the agent’s decision-making process and evidence weighting.
Discussion and Limitations
Our AIFo framework potentially represents a paradigm shift in AI-generated image detection by emulating human forensic reasoning through multi-agent collaboration. The framework’s training-free nature and cross-model generalizability address key limitations of existing detection methods, offering a more sustainable and adaptable solution for the rapidly evolving landscape of generative AI.
Despite the promising results, our approach has several important limitations that warrant careful consideration.
Scalability and Computational Efficiency. While our multi-agent approach achieves high accuracy, it comes with increased computational overhead compared to single-model solutions. The sequential and parallel execution of multiple forensic tools, combined with LLM-based reasoning, results in higher latency and resource consumption. For large-scale deployment scenarios, optimization strategies such as tool prioritization, caching mechanisms, and selective tool activation based on image characteristics could help balance accuracy and efficiency.
Dependency on External Services. The framework’s effectiveness is inherently tied to the availability and reliability of external services, particularly reverse image search APIs and metadata extraction tools. Changes in API policies, service outages, or modifications to search algorithms could impact the framework’s performance. This dependency creates potential points of failure that are beyond the system’s direct control.
Adversarial Metadata Manipulation. Our framework lies in its reliance on image EXIF metadata as a key source of forensic evidence. However, adversaries could potentially manipulate image metadata to mislead the detection system. For instance, attackers could inject fake EXIF data mimicking legitimate camera parameters into AI-generated images. Such metadata spoofing attacks could compromise the reliability of our metadata extraction tool, which currently shows high reliability rates in our evaluation.
Related Work
Fake Image Detection
Early approaches to AI-generated image detection mainly rely on training image classifiers using machine learning, often using datasets produced by specific generative models. Over time, research has shifted towards more generalizable and robust detection strategies, including leveraging the visual capabilities of large multimodal models (LMMs) such as CLIP [42]. CNNSpot [53] is one of the first works to propose a “universal” detector capable of distinguishing real images from CNN-generated ones without being dependent on a particular network architecture or training dataset. De-Fake [46] extends the detection pipeline by incorporating both the image content and generated prompts using CLIP to train a hybrid classifier. DIRE [54] and ZeroFake [47] exploit intrinsic differences between real and fake images revealed during the diffusion model reconstruction process to build detection models with improved generalization. PatchCraft [59] focuses on local texture patches, identifying subtle artifacts left by generative models in fine-grained regions.
Recent works utilize VLMs as fake image detectors, such as AntifakePrompt [18], the method proposed by Jia et al. [29], and the approach by Ji et al. [28]. AIGI-Holmes [60] proposes a complete framework to train a VLM for explainable and generalizable detection, aiming to produce human-verifiable justifications. The work by Yu et al. [57] develops a framework to enhance generalization and explainability by using a knowledge-guided detector and a forgery-aware prompt learner. FakeBench [33] and DFBench [51] introduced a large-scale benchmark to rigorously test the detection performance of LMMs against a wide range of modern generative models.
Despite these progresses, existing methods exhibit inherent limitations. These approaches predominantly rely on internal, pixel-level visual features, while overlooking complementary external sources of evidence that can be critical for reliable forensic analysis. In contrast to standalone detector architectures, our work adopts an agent-based framework that emulates the investigative workflow of human forensic experts. The proposed AI agent leverages a diverse set of external tools to systematically collect, integrate, and reason over different evidence, which allows for a more robust, context-aware, and explainable framework for AI-generated image detection.
LLM-Based Multi-Agent Frameworks
Recent advances in large language models (LLMs) have catalyzed the development of agentic frameworks that simulate complex human-like workflows by coordinating multiple specialized agents. ReAct [56] introduces a framework that switches between reasoning and acting within language models. Other works such as MetaGPT [24] simulate various roles in a software company and build a multi-agent software development framework. In the security domain, several studies have developed multi-agent frameworks tailored to diverse security-related downstream tasks, including adversarial defense [58], harmful content detection [36], bias identification in generative models [52], fake news verification [32], and comparative evaluation of misinformation detection strategies [27]. However, existing approaches do not specifically target zero-shot detection tasks nor address the unique challenges of AI-generated image detection. To the best of our knowledge, our work is the first to develop a training-free, zero-shot multi-agent framework specifically designed for AI-generated image detection. Combining the strengths of existing image detection methods and LLM agents, we present a novel and practical solution for addressing the evolving challenges in AI-generated image forensics.
Conclusion
In this work, we introduce AIFo, a novel multi-agent framework that advances AI-generated image detection by emulating human forensic reasoning. Unlike conventional methods, AIFo integrates diverse forensic tools and uses multi-agent collaboration to synthesize evidence. Our evaluation on a comprehensive benchmark spanning both lab and real-world settings shows that AIFo achieves 97.05% accuracy, substantially outperforming traditional classifiers and state-of-the-art VLMs like GPT-4o.
The key contributions include a training-free, agent-based paradigm that ensures generalizability through procedural reasoning, a finding that multi-source evidence integration and structured debate mechanisms significantly enhance both accuracy and interpretability, and the establishment of a comprehensive benchmark that enables evaluation under real-world conditions. While our framework represents a significant advancement in AI-generated image detection, it also faces some limitations including computational overhead and dependency on external services.
As generative AI technologies continue to evolve rapidly, the need for robust, interpretable, and adaptable detection systems becomes increasingly critical. Our agent-based approach offers a sustainable foundation for addressing these challenges, providing a framework that can generalize alongside advancing generative models while maintaining the transparency and reliability essential for real-world deployment in security-critical applications.
Ethical Considerations
In this study, we adopted a stakeholder-oriented perspective to examine the ethical dimensions of our work. For the research team, the development and validation of our new detection framework contributed to advancing technical expertise and academic reputation. For the general public, the framework offers a practical tool to mitigate the spread of misinformation by improving the detection of AI-generated images, thereby safeguarding individuals from being misled. Companies such as social media platforms and news organizations may also benefit by employing the framework to verify content authenticity and maintain the credibility of their services. Our research is guided by several core ethical principles. First, the principle of beneficence is reflected in our aim to protect society from the harmful consequences of misinformation. Second, respect for persons is ensured by using only publicly available datasets that do not involve personal or sensitive information. Third, the principle of justice informed our effort to design a framework whose outcomes can be applied broadly across different social groups, thereby promoting fair access to reliable information. We think that this research provides substantial value in promoting information authenticity and strengthening public trust. We are therefore confident that the study is ethically sound and makes a meaningful contribution to the ongoing development of AI-generated image detection.
Appendix A Prompts Used in Our Framework
This appendix documents the exact prompts used in our experiments to facilitate reproducibility. We group prompts according to their roles in the multi-agent framework.
Evidence Gatherer Agent Prompt
The Evidence Gatherer Agent collects cross-source forensic signals. Table 8 is the prompt template used to instruct the LLM:
You are an AI Image Forensics Expert. Your task is to determine whether the input image is AI-generated or real using the available forensic tools.
- A real image refers to images created by humans, including photographs captured by cameras, photos that have been edited with software such as Photoshop, or human artistic creations such as hand-drawn sketches and paintings.
- An AI-generated image refers to images that is fully or partially generated by AI models.
Available Tools:
- reverse_search: Perform a reverse image search to find exact matches or similar appearances online.
- extract_image_metadata: Inspect technical EXIF metadata for authenticity cues.
- vlm_analysis: Obtain expert-level visual analysis of the image content.
- pre-trained_classifiers: Apply dedicated AI-generated image detection models.
Your role is to systematically invoke these tools as needed and collect evidence that will later be assessed to determine the authenticity of the input image.
Table 8: Prompt template for the Evidence Gatherer Agent.
Reasoning Agent Prompt
The Reasoning Agent first assesses whether the evidence is sufficient and consistent to support a decision. If so, it synthesizes all sources to produce a final binary judgment with a explanation that evaluates source reliability. Table 9 and Table 10 are the prompt templates used to instruct the LLM:
You are an AI Image Forensics Expert. Your task is to determine if the following evidence collected from multiple tools is sufficient and consistent enough to make a final judgment.
{tool_results}
Answer ’True’ if the evidence is both sufficient and consistent enough to confidently reach a final decision and ’False’ if the evidence is incomplete, ambiguous, or contains major conflicts that require further debate and analysis.
Table 9: First prompt template for the Reasoning Agent.
You are an AI Image Forensics Expert. Your task is to determine whether the image ia ai-generated or a real image.
- A real image refers to images created by humans, including photographs captured by cameras, photos that have been edited with software such as Photoshop, or human artistic creations such as hand-drawn sketches and paintings.
- An AI-generated image refers to images that is fully or partially generated by AI models.
Please make a final judgment based on the following evidence collected from multiple tools:
{tool_results}
Critically evaluate each evidence source and its reliability.
Required output format:
1. is_ai_generated: boolean (True if AI-generated, False if real image)
2. analysis_details: A detailed analysis explaining your decision
Table 10: Second prompt template for the Reasoning Agent.
Debate Agents Prompt
The Debate Agents engage in a structured debate to resolve conflicts and ambiguities in the evidence. Table 11 is the prompt template used to instruct the Pro-Agent LLM:
You are an AI Image Forensics Expert. Your goal is to correctly classify an image as either AI-generated or real.
Your analysis must be based on the evidence provided in the tool results below.
Tool Results:
{tool_results}
#First Round Only:
You are arguing in favor of the image being AI-generated.
Scrutinize the tool results for any artifacts, inconsistencies, or patterns typical of AI generation. Present your findings as a concise, bullet-pointed list. Focus on the strongest pieces of evidence that support your assigned perspective.
#Subsequent Rounds Only:
Review the other expert’s points from the previous round and re-evaluate your own position.
- Acknowledge any valid points they made.
- Re-examine the tool results to see if their perspective reveals something you missed.
- Refine or strengthen your analysis based on this new information. Your updated analysis should be more nuanced.
You are arguing in favor of the image being AI-generated.
The other expert’s (arguing for "Real") points:
{negative_history}
Provide your updated, refined analysis as a concise bullet-pointed list.
Table 11: Prompt template for the Pro-Agent.
Table 12 is the prompt template used to instruct the Con-Agent LLM:
You are an AI Image Forensics Expert. Your goal is to correctly classify an image as either AI-generated or real.
Your analysis must be based on the evidence provided in the tool results below.
Tool Results:
{tool_results}
First Round Only:
You are arguing in favor of the image being authentic (real).
Look for signs of naturalness, photographic properties, and details that are hard for AI to replicate, based on the tool results.
Subsequent Rounds Only:
Review the other expert’s points from the previous round and re-evaluate your own position.
- Acknowledge any valid points they made.
- Re-examine the tool results to see if their perspective reveals something you missed.
- Refine or strengthen your analysis based on this new information. Your updated analysis should be more nuanced.
You are arguing in favor of the image being authentic (real).
The other expert’s (arguing for "AI-generated") points:
{positive_history}
Provide your updated, refined analysis as a concise bullet-pointed list.
Table 12: Prompt template for the Con-Agent.
Judge Agent Prompt
The Judge Agent is tasked with overseeing the debate process and synthesizing the final decision based on both the tool-derived evidence and the debate history. The Judge also evaluates the sufficiency of each debate round and can decide to terminate the debate early if the arguments are deemed sufficient. Table 13 and Table 14 are the prompt templates used to instruct the LLM:
As an impartial judge, review the debate history so far.
Your task is NOT to make the final decision, but to determine if the debate is sufficient to support a final decision.
Arguments for ’AI-generated’:
{positive_args}
Arguments for ’Authentic Image’:
{negative_args}
Your Decision Criteria:
1. If one side’s evidence is strong and the other’s is weak or has been effectively countered, the information is likely sufficient.
2. If both sides have presented compelling but conflicting evidence that has not yet been reconciled, more analysis is needed.
3. If the discussion become repetitive, further rounds are unlikely to be productive.
Based on these criteria, decide if you have enough information to make a high-confidence final judgment.
Answer ’True’ if sufficient, ’False’ if more debate and analysis would be helpful.
Table 13: First prompt template for the Judge Agent.
You are an AI Image Forensics Judge. Your role is to synthesize all available information and deliver a definitive, well-reasoned judgment on whether the image is AI-generated or real.
- A real image refers to images created by humans, including photographs captured by cameras, photos that have been edited with software such as Photoshop, or human artistic creations such as hand-drawn sketches and paintings.
- An AI-generated image refers to images that is fully or partially generated by AI models.
Raw Evidence from tools: tool_results Arguments for ’AI-generated’:
{positive_args}
Arguments for ’Authentic Image’:
{negative_args}
Your analysis must be a comprehensive synthesis. Follow these steps in your reasoning:
1. Weigh the Evidence: Identify the most compelling piece of evidence from EACH side.
2. Resolve the Core Conflict: Directly address the central disagreement.
3. State Your Final Conclusion: Based on your analysis, provide a clear final verdict.
Required output format:
1. is_ai_generated: boolean (True if AI-generated, False if real image)
2. analysis_details: A detailed analysis explaining your decision
Format the response as a structured object.
Table 14: Second prompt template for the Judge Agent.
VLM Analysis Tool Prompt
The VLM Analysis Tool utilizes vision-language models to conduct in-depth visual analysis of images. The prompt used to guide the VLM Analysis Tool is detailed in Table 15, ensuring the model focuses on key visual characteristics and provides comprehensive evidence to support its classification.
As a professional AI image detector, please analyze this image carefully:
1. Determine if this is an AI-generated image or a real image.
- Real images include images that are created by humans, including photographs captured by cameras, photos that have been edited with software such as Photoshop, or human artistic creations such as hand-drawn sketches and paintings.
- AI-generated images include images that are fully or partially generated by AI models.
2. If you determine it’s an AI-generated image, please specifically identify and list the visual artifacts or characteristics that indicate AI generation, such as:
- Unnatural textures or patterns
- Inconsistent lighting or shadows
- Anatomical errors in humans or animals
- Unusual distortions or blending of elements
- Text or writing abnormalities
- Symmetry issues or repeating patterns
- Unusual backgrounds or contextual inconsistencies
3. If you determine it’s a real image, explain what characteristics support this conclusion.
4. Provide your final classification with confidence level (high, medium, or low).
Table 15: Prompt template for the VLM Analysis Tool.
Appendix B Metadata Analysis Tool Key Fields
Table 16 provides the exact key fields and prefixes used in the metadata analysis tool to identify authenticity markers in images.
Table 16: Metadata fields and prefixes considered in the analysis tool.
Category Field / Prefix Description
Exact Key Fields
XMP:CreatorTool Creator tool Software used to generate or edit the image.
EXIF:Software Software tag Image editing or generation software information.
EXIF:UserComment User comment Arbitrary comments added to the image metadata.
File:Comment File comment Comments embedded directly in the file container.
XMP:Description Description Textual description of the image.
XMP:Title Title Title field embedded in XMP metadata.
XMP:Rights Rights Usage rights or copyright information.
XMP:Source Source Original source reference of the image.
EXIF:Make Camera make Manufacturer of the recording equipment.
EXIF:Model Camera model Camera model used for the photo.
EXIF:LensModel Lens model Lens information recorded by the camera.
EXIF:LensInfo Lens info Technical specifications of the lens.
EXIF:LensSerialNumber Lens serial number Unique identifier of the lens.
EXIF:ExposureTime Exposure time Shutter exposure duration.
EXIF:FNumber F-number Aperture size of the lens.
EXIF:ISO ISO Sensitivity setting of the camera.
EXIF:FocalLength Focal length Lens focal length value.
EXIF:SerialNumber Camera serial number Unique identifier of the camera.
EXIF:GPSLatitude GPS latitude Geographic latitude of capture.
EXIF:GPSLongitude GPS longitude Geographic longitude of capture.
EXIF:GPSTimeStamp GPS timestamp Time recorded by GPS.
EXIF:DateTimeOriginal Original datetime Original capture time of the image.
EXIF:CreateDate Creation date File creation date.
Composite:GPSPosition GPS position Combined GPS coordinates.
Composite:Aperture Aperture Derived aperture value.
Composite:ShutterSpeed Shutter speed Derived shutter speed.
Composite:LensID Lens ID Identifier for the lens model.
ICC_Profile:ProfileDescription ICC profile description Description of the color profile.
ICC_Profile:ProfileCopyright ICC profile copyright Copyright information for the ICC profile.
IPTC:DocumentNotes Document notes Notes in IPTC metadata.
IPTC:ApplicationRecordVersion Record version Version of the IPTC application record.
Key Field Prefixes
MakerNotes: Camera-specific notes Manufacturer-specific EXIF metadata.
JUMBF: JUMBF metadata Metadata block for embedding auxiliary information.
MPF: Multi-picture format Metadata for multi-frame images.
Appendix C Detailed Accuracy performance
Table 17 provides a detailed breakdown of the accuracy performance of various methods across different image sources, including both in-the-lab and in-the-wild scenarios.
Table 17: Detailed Accuracy performance of different methods on each image sources.
In-the-Lab In-the-Wild
Real Images AI Images Real Images AI Images
Method Flickr30k ImageNet DIV2k GenImage FakeBench Holopix50k Flickr W. Commons Lexica Nightcafe Civitai
CNNSpot [53] 0.9980 0.9980 1.0000 0.0475 0.1857 1.0000 1.0000 0.9980 0.0000 0.0020 0.0000
PatchCraft [59] 0.9900 0.9140 0.7680 0.8525 0.5986 0.8600 0.9040 0.5960 0.0140 0.2360 0.3360
DE-FAKE [46] 0.8980 0.7280 0.3480 0.7225 0.6443 0.7180 0.5700 0.4920 0.9840 0.9380 0.8360
GPT-4.1 [8] 1.0000 0.9920 0.9880 0.8350 0.9086 0.9980 0.9860 0.9920 0.7840 0.9900 0.9500
GPT-4o [39] 1.0000 0.9860 0.9960 0.8850 0.9471 0.9980 0.9820 0.9940 0.7520 0.9900 0.9420
AIFo (ours) 1.0000 0.9840 0.9920 0.9475 0.9657 0.9960 0.9880 0.9940 0.8420 0.9880 0.9940
Appendix D AI Model Sources
Table 18 provides an overview of the AI models used for generating images in our benchmark’s AI-sourced datasets.
Table 18: Overview of the AI models and platforms used for generating images in our benchmark’s AI-sourced datasets. The table is categorized by the In-the-Lab and In-the-Wild settings.
Dataset Source AI Models and Platforms Used for Generation
In-the-Lab AI Image Sources
GenImage [61] BigGAN, GLIDE, VQDM, ADM, Midjourney, Wukong, and Stable Diffusion (v1.4, v1.5).
FakeBench [33] ProGAN, StyleGANs, CogView2, FuseDream, VQDM, GLIDE, Midjourney, Stable Diffusion, DALL·E 2, and DALL·E 3.
In-the-Wild AI Image Sources
Lexica [9] Lexica Aperture Series (v3.5, v4, v5, Max).
Nightcafe [10] DALL·E 2, DALL·E 3, Stable Diffusion, and various other community fine-tuned models.
Civitai [5] A vast collection of community fine-tuned models, predominantly based on Stable Diffusion (including SDXL variants) series.

### v2

@@
 Abstract
-The rapid evolution of AI-generated images poses unprecedented challenges to information integrity and media authenticity.
-Existing detection approaches suffer from fundamental limitations: traditional classifiers lack interpretability and fail to generalize across evolving generative models, while vision-language models (VLMs), despite their promise, remain constrained to single-shot analysis and pixel-level reasoning.
-To address these challenges, we introduce AIFo (Agent-based Image Forensics), a novel training-free framework that emulates human forensic investigation through multi-agent collaboration.
-Unlike conventional methods, our framework employs a set of forensic tools, including reverse image search, metadata extraction, pre-trained classifiers, and VLM analysis, coordinated by specialized LLM-based agents that collect, synthesize, and reason over cross-source evidence.
-When evidence is conflicting or insufficient, a structured multi-agent debate mechanism allows agents to exchange arguments and reach a reliable conclusion.
-Furthermore, we enhance the framework with a memory-augmented reasoning module that learns from historical cases to improve future detection accuracy.
-Our comprehensive evaluation spans 6,000 images across both controlled laboratory settings and challenging real-world scenarios, including images from modern generative platforms and diverse online sources.
-AIFo achieves 97.05% accuracy, substantially outperforming traditional classifiers and state-of-the-art VLMs.
-These results demonstrate that agent-based procedural reasoning offers a new paradigm for more robust, interpretable, and adaptable AI-generated image detection.
+The rapid evolution of AI-generated images poses growing challenges to information integrity and media authenticity.
+Existing detection approaches face limitations in robustness, interpretability, and generalization across diverse generative models, particularly when relying on a single source of visual evidence.
+We introduce AIFo (Agent-based Image Forensics), a training-free framework that formulates AI-generated image detection as a multi-stage forensic analysis process through multi-agent collaboration.
+The framework integrates a set of forensic tools, including reverse image search, metadata extraction, pre-trained classifiers, and vision-language model analysis, and resolves insufficient or conflicting evidence through a structured multi-agent debate mechanism.
+An optional memory-augmented module further enables the framework to incorporate information from historical cases.
+We evaluate AIFo on a benchmark of 6,000 images spanning controlled laboratory settings and challenging real-world scenarios, where it achieves 97.05% accuracy and consistently outperforms traditional classifiers and strong vision-language model baselines.
+These findings demonstrate the effectiveness of agent-based procedural reasoning for AI-generated image detection.
 Introduction
-In recent years, image generative models such as GLIDE [37], Imagen [45], DALL·E 2 [43], and Stable Diffusion [44], have advanced rapidly.
-They can synthesize photorealistic images from natural language in seconds [23, 44, 45].
-However, the realism of AI-generated images has raised serious societal concerns.
-Because AI-generated images can now easily fool human observers, malicious actors are increasingly leveraging them to spread disinformation and impersonate individuals [48].
-For example, during the 2024 U.S.
-presidential election cycle, sophisticated deepfakes (i.e., AI-generated images or videos that convincingly fabricate real people or events) have appeared in campaign ads and on social media, which potentially manipulate public opinions and disrupt the voting behaviors [15].
-Beyond elections, AI-generated images also introduce broader risks [48, 50], including misinformation and privacy infringement.
+Figure 1: High-level overview of our proposed AIFo.
+Recent years have witnessed rapid advancements in image generative models, ranging from early works such as GLIDE [NDRSMMSC21], Imagen [SCSLWDGAMLSHFN22], DALL·E 2 [RDNCC22], and Stable Diffusion [RBLEO22] to more recent foundation models including HunyuanImage [CCCCCDDGGGo25] and Qwen-Image [WLZLGYYBXCCTZWYYCLLZMWNCCPQWWYWFXWZZWCL25].
+They can synthesize photorealistic images from natural language in seconds.
+However, the realism of AI-generated images has raised serious societal concerns, as they can be used for disinformation, impersonation, and privacy infringement, undermining public trust [SWAZ23, V20].
 In response to these risks, substantial research has been devoted to the detection of AI-generated images.
-Current methodologies can be generally classified into two main categories: traditional machine learning classifiers and advanced approaches leveraging large vision language models (VLMs).
-Traditional machine learning classifiers typically rely on training convolutional neural networks (CNNs) or transformer-based models to distinguish between real and fake images [53, 59, 54, 19, 17].
-Early studies reveal that AI-generated images tend to exhibit shared low-level artifacts, allowing detectors trained on labeled images to identify them [53, 46].
-For example, DE-FAKE [46] trains a set of classifiers on AI-generated and real images to learn AI-specific artifacts.
-However, these works often depend on a limited number of training datasets with fake images produced by only a few specific generative models, which makes them prone to generalizability issues: they perform well on seen data but struggle to generalize to unseen images from new generative models [60, 51].
-They also often lack explainability: most models act as black boxes, producing binary outputs without offering human-interpretable justifications [33, 60].
-More recently, vision language models (VLMs) have shown promise for more generalizable and explainable detection [60, 57, 33, 28].
-Due to the large-scale pre-training, VLMs can be transferred to image detection tasks in a zero-shot or few-shot manner [60, 57, 51], without relying on specialized labeled datasets.
-Beyond identifying pixel-level artifacts like traditional classifiers, VLMs can also apply semantic and world knowledge in the detection process, e.g., a photo of a flying cat is AI-generated because a flying cat cannot exist in the physical world.
-In addition, they can provide human-interpretable justifications through prompt engineering for explainable detection [57, 28, 51].
-Despite these advancements, both traditional and VLM-based approaches suffer from fundamental limitations compared to human forensic experts.
-First, they rely heavily on pixel-level image features as the main detection evidence.
-By contrast, human experts can not only employ off-the-shelf classifiers and VLMs for pixel-level analysis, but also actively seek evidence beyond the images themselves, e.g., Exchangeable Image File Format (EXIF) metadata that contains important information about the image or online contextual information to make a comprehensive decision.
-Second, both approaches treat image detection as a single-shot classification task, using a fixed model or pipeline.
-Human experts, on the other hand, approach image detection as a dynamic reasoning process: they flexibly employ different tools and iteratively refine their judgments based on the gathered evidence.
-Finally, human experts can improve over time through accumulated experience, enabling them to adapt to new generative models and real-world scenarios, whereas fixed models remain static unless frequently fine-tuned on new datasets.
-Motivated by this, we aim to explore a new paradigm that combines the advantages of existing classifiers and VLMs with the capabilities of human experts for AI-generated image detection.
-Our Work.
-Our approach fundamentally differs from conventional methods based on static classifiers or single VLM.
-Rather than developing another standalone image classifier, we design a cognitive AI agent system that automates the entire human-like forensic workflow, simulating reasoning and decision-making in the detection of AI-generated images.
-We present AIFo (Agent-based Image Forensics), an LLM-based multi-agent framework that integrates functionalities such as visual recognition, semantic understanding, and provenance analysis.
-AIFo consists of a forensic Toolbox, an Evidence Gatherer, a Reasoning Agent, and a Debate module.
-Given a test image, the Evidence Gatherer first applies tools from the Toolbox and aggregates the resulting evidence.
-The Reasoning Agent then evaluates the quality of the collected evidence, i.e., whether it is sufficient and consistent enough to support a reliable judgment.
-If so, the Reasoning Agent produces a final decision and explains why the image is classified as AI-generated or real.
-Otherwise, we introduce a Debate module [21, 35, 36, 34, 49] to handle cases where the evidence is insufficient or conflicting.
-In this module, two Debate Agents exchange arguments over multiple rounds, each adopting an opposing stance (for or against the claim that the image is AI-generated), while a Judge Agent oversees the debate process and produces the final judgment.
-Additionally, we demonstrate the potential of our framework to continuously and progressively enhance its detection effectiveness by learning from historical testing data.
-By incorporating a memory module that stores all testing history, the framework behaves like a human expert, accumulating experience and improving performance over time.
-Overall, such a cognitive AI agent framework offers a new paradigm for AI-generated image detection.
-Main Findings.
-We evaluate AIFo on a dataset of 6,000 AI-generated and real images, comprising 3,000 samples from five established benchmarks (e.g., Flickr30K [40], GenImage [61], and FakeBench [33]) and 3,000 in-the-wild images collected from six online platforms.
-AIFo is benchmarked against a range of baselines methods, including traditional classifiers such as CNNSpot [53], DE-FAKE [46], and state-of-the-art VLMs (e.g., GPT-4.1 [8] and GPT-4o [39]).
-To ensure a fair comparison, we disable the memory module and prevent our framework from learning during the main evaluation.
-Even under these conditions, AIFo achieves the best overall performance, reaching 0.9705 accuracy and surpassing GPT-4o (0.9483), GPT-4.1 (0.9416), and other baselines.
-It also maintains higher robustness (0.9047–0.9690) than GPT-4o (0.8818–0.9462) under three types of perturbations.
-To demonstrate the effectiveness of the memory module, we conduct a case study on 50 images that are misclassified in the main evaluation.
-The results show that when the memory module stores similar historical cases, approximately 40% of these errors are successfully corrected.
-The main contributions of our work are as follows:
-  • Human-Like Procedural Reasoning.
-We present the first agentic framework (AIFo) that simulates human-like procedural reasoning for AI-generated image detection.
-Our framework combines the strengths of conventional classifiers and VLMs, employing them as complementary tools, and incorporates human-like reasoning processes such as integrating multiple sources of evidence, debating conflicting evidence, and improving through accumulated experience.
-This highlights a transition from static classification toward a dynamic reasoning process in AI-generated image detection.
-  • Comprehensive Evaluation.
-We contribute a valuable benchmark dataset covering 6000 images from both existing benchmark datasets and in-the-wild images from internet platforms.
-This dataset serves as the foundation for rigorous evaluation, through which we demonstrate the superior performance, generalizability, and robustness of our AIFo compared to state-of-the-art baselines.
-  • Training-Free Core and Cross-Model Generalizability.
-Unlike conventional detectors that require large-scale training data and frequent updates, the core of AIFo is training-free and designed to generalize across evolving generative models.
-It leverages generalizable forensic evidence rather than model-specific features, enabling robust detection of AI-generated images.
-An optional memory module further enhances performance over time by accumulating historical cases.
-Threat Model
-Detector’s Goals
-Distinguishing Between AI-Generated and Real Images.
-The detector aims to determine whether a given image is AI-generated or a real image.
-In this work, AI-generated images include those produced by generative models like text-to-image diffusion models, as well as AI-edited images, such as cartoons generated from real photographs and images replaced with AI-generated backgrounds.
-Real images refer to those captured by physical cameras or manually created/drawn by human artists.
-Notably, if a real image has been edited or post-processed by humans (without the aid of AI models), e.g., by applying filters, adjusting color contrast, or increasing brightness, we still consider it a real image.
-Enhanced Interpretability.
-Unlike conventional binary detectors, our detector aims not only to provide a decision but also to offer human-interpretable justifications.
-For example, if our detector identifies an image as an AI-generated image, it should also point out evidence that supports this decision, such as unreasonable object placement (semantic-level evidence), distorted lighting in the image (pixel-level evidence), editing history (metadata-level evidence), AI watermarks (source-level evidence), etc.
-Providing such explanations helps users understand, verify, and trust the detector’s decisions.
-Cross-Model Generalizability.
-Given the rapid evolution of text-to-image generative models, new architectures and generated images are continually emerging, which may degrade the performance of detectors trained on previous generative models.
-Therefore, the detector should be robust to various model iterations and new architectures, and capable of leveraging generalizable evidence to assess image provenance, rather than relying solely on model-specific features.
-Training-Free Design.
-Furthermore, traditional AI image detection frameworks typically depend on collecting large-scale training datasets of AI-generated images.
-As generative models evolve, these frameworks require continual updates to their training data, incurring significant maintenance costs.
-In contrast, the core of our detector is designed to be training-free: it does not require any dedicated training set, and can be directly applied to detect AI-generated images from any model, including those unseen during development.
-Detector’s Capabilities
-Our detector is assumed to have access to the following resources and conditions:
-  • Image-File Access. The detector can access the image file itself, including its metadata if available (e.g., camera details, timestamps, and editing history).
-  • External Forensic Tools.
-The detector is permitted to utilize a suite of external forensic tools to analyze the image and gather supporting evidence.
-This includes pretrained AI detection models, reverse image search engines, and vision-language models.
+Current methodologies can be generally classified into two main categories: traditional machine learning classifiers and approaches leveraging large vision-language models (VLMs).
+Traditional machine learning classifiers typically rely on training convolutional neural networks (CNNs) or transformer-based models to distinguish between real and fake images [WWZOE20, ZXLQZ24, WBZWHCL23, CPNV24, BLYXLHC23].
+Early studies reveal that AI-generated images tend to exhibit shared low-level artifacts, allowing detectors trained on labeled images to identify them [WWZOE20, SLYZ22].
+For example, DE-FAKE [SLYZ22] trains a set of classifiers on AI-generated and real images to learn AI-specific artifacts.
+While effective under controlled settings, such approaches often struggle to generalize to unseen generative models and provide limited insight into how individual cues contribute to the final decision [ZLWSJYDSWJ25, WDWJYZZQXZM25, LLWLWRL24].
+More recently, VLMs have shown promise for more generalizable and explainable detection [ZLWSJYDSWJ25, YFGFXC25, LLWLWRL24, JHZCLZWZZ25].
+Due to the large-scale pre-training, VLMs can be transferred to image detection tasks in a zero-shot or few-shot manner [ZLWSJYDSWJ25, YFGFXC25, WDWJYZZQXZM25], without relying on specialized labeled datasets.
+Moreover, through prompt engineering, VLMs are able to produce interpretable justifications alongside their predictions [YFGFXC25, JHZCLZWZZ25, WDWJYZZQXZM25].
+Despite these advancements, both traditional and VLM-based methods remain limited compared to human forensic experts.
+Human experts rarely rely on a single visual cue, instead integrating diverse evidence sources and iteratively refining their conclusions as new information emerges.
+By contrast, most existing methods perform fixed classification and lack explicit mechanisms to assess evidence sufficiency, resolve conflicting signals, or accumulate experience over time.
+Motivated by this, we seek a new paradigm that integrates the strengths of classifiers and VLMs with the procedural reasoning capabilities exhibited by human experts for AI-generated image detection.
+Instead of developing another image classifier, we argue that AI-generated image detection should be viewed as an evidence-driven decision-making problem.
+Therefore, we adopt a procedural reasoning formulation inspired by human forensic workflows, in which a verdict is reached through multiple stages of reasoning rather than a single forward pass.
+We present AIFo (Agent-based Image Forensics), a training-free, multi-agent framework that coordinates multiple forensic tools through specialized agent roles.
+As shown in Figure 1, AIFo structures the detection process into distinct stages, including evidence collection, evidence assessment, and an optional debate stage.
+An Evidence Gatherer Agent invokes tools from a forensic toolbox and aggregates their outputs, while a Reasoning Agent evaluates whether the collected evidence is sufficient and consistent to support a reliable decision.
+When evidence is ambiguous or contradictory, a debate mechanism is activated, allowing opposing hypotheses to be examined under the supervision of a Judge Agent.
+We further introduce a memory module that stores historical cases, enabling the framework to accumulate experience and improve performance over time.
+We evaluate AIFo on a dataset of 6,000 AI-generated and real images, comprising 3,000 samples from five established benchmarks (e.g., Flickr30K [PWCCHL15], GenImage [ZCYHLLTHHW23], FakeBench [LLWLWRL24], among others) and 3,000 real-world images from six online platforms.
+Across these settings, AIFo consistently outperforms traditional classifiers such as CNNSpot [WWZOE20] and DE-FAKE [SLYZ22], as well as strong VLM baselines (e.g., GPT series [O23]), achieving higher accuracy and robustness under common perturbations.
+The main contributions of our work are: (1) We reformulate AI-generated image detection as an evidence-driven forensic decision problem, highlighting the role of evidence sufficiency, consistency, and conflict resolution under uncertainty.
+Building on this, we present an agent-based procedural reasoning framework that implements key stages of human forensic analysis, including evidence aggregation, evidence assessment, and structured conflict resolution.
+(2) Our proposed framework AIFo is a training-free system that integrates conventional classifiers and VLMs as complementary evidence sources rather than standalone detectors, with a modular design that enables robust generalization to evolving generative models.
+(3) We conduct extensive empirical evaluations on a benchmark of 6,000 images combining public datasets and real-world samples, and demonstrate the superior performance, generalizability, and robustness of AIFo over state-of-the-art baselines.
+Related Work
+Fake Image Detection.
+Early approaches to AI-generated image detection mainly rely on training image classifiers using machine learning, often using datasets produced by specific generative models.
+Over time, research has shifted towards more generalizable and robust detection strategies, including leveraging the visual capabilities of large multimodal models (LMMs) such as CLIP [RKHRGASAMCKS21].
+CNNSpot [WWZOE20] proposed one of the first universal detectors independent of generator architecture or dataset.
+De-Fake [SLYZ22] combined image content and textual prompts using CLIP-based hybrid training.
+DIRE [WBZWHCL23] and ZeroFake [STLBZ24] exploit intrinsic differences between real and fake images revealed during the diffusion model reconstruction process to build detection models with improved generalization.
+PatchCraft [ZXLQZ24] focuses on local texture patches, identifying subtle artifacts left by generative models in fine-grained regions.
+AIDE [YLCHJHX24] detects AI-generated images by selecting highest and lowest frequency patches through a mixture-of-experts architecture.
+Recent research leverages VLMs for detection, such as AntifakePrompt [CYCY23], Jia et al.
+[JLZCYJHLWL24], and Ji et al.
+[JHZCLZWZZ25].
+AIGI-Holmes [ZLWSJYDSWJ25] proposes a complete framework to train a VLM for explainable and generalizable detection, aiming to produce human-verifiable justifications.
+The work by Yu et al.
+[YFGFXC25] develops a framework to enhance generalization and explainability by using a knowledge-guided detector and a forgery-aware prompt learner.
+FakeBench [LLWLWRL24] and DFBench [WDWJYZZQXZM25] introduced a large-scale benchmark to rigorously test the detection performance of LMMs against a wide range of modern generative models.
+Despite these advances, existing methods still mainly rely on internal, pixel-level visual features, overlooking complementary external evidence critical to forensic reasoning.
+In contrast, our agent-based framework emulates human investigative workflows, leverages external tools and reason over different evidence, which allows for a robust and explainable framework for AI-generated image detection.
+LLM-Based Multi-Agent Frameworks.
+Recent advances in LLMs have catalyzed the development of agentic frameworks that simulate complex human-like workflows by coordinating multiple specialized agents.
+ReAct [YZYDSNC23] introduces a framework that switches between reasoning and acting within LLMs.
+Other works such as MetaGPT [HZCZCWZWYLZRXWS24] simulate various roles in a software company and build a multi-agent software development framework.
+Moreover, multi-agent frameworks have also been applied to adversarial defense [ZWZWW24], harmful content detection [LFLWD25], bias identification in generative models [WBYXCMZCW23], fake news verification [LZM24], and misinformation evaluation [HYYX25].
+However, existing approaches primarily emphasize predictive accuracy and explainability, while leaving evidence aggregation and conflict resolution largely implicit.
+We therefore present the first training-free multi-agent framework tailored to this task, combining the strengths of existing detectors and LLM agents for adaptive, explainable AI image forensics.
 The AIFo Framework
 Design Rationale
-Unlike existing approaches that rely solely on pre-trained classifiers or VLMs, our main design principle is to emulate the procedural reasoning processes employed by human forensic experts [22, 38], while still leveraging the strengths of existing approaches.
-To this end, we leverage LLMs as autonomous agents to develop an agentic framework that incorporates the key capabilities of human forensic experts.
-First, the framework is designed to employ and coordinate a suite of forensic tools, such as pre-trained classifiers, metadata extraction, reverse image search, and VLM-based analysis.
-Second, rather than relying on a single fixed source of evidence, the framework should be able to cross-check evidence from multiple tools before forming the final decision.
-Next, when the collected evidence indicates conflicting signals, the framework is expected to resolve these conflicts through human-like reasoning activities, such as assessing the reliability of evidence sources and engaging in structured debate among the expert team.
-The core framework is training-free, building on top of pre-trained classifiers and state-of-the-art VLMs as forensic tools.
-At the same time, an optional memory module allows the framework to accumulate historical cases and progressively improve performance.
-Altogether, we aim to build a training-free, interpretable, and robust framework for AI-generated image detection, which not only leverages existing detection approaches but also possesses key capabilities of human forensic experts, e.g., integrating various evidence, evaluating their reliability, and resolving conflicts among them.
+Instead of treating AI-generated image detection as a single-shot visual classification problem, we formulate it as an evidence-driven reasoning process that operates under heterogeneous, incomplete, and sometimes conflicting evidence, emulating the procedural reasoning of human forensic experts [FH24, NLBFPCL24].
+Our framework follows a multi-stage reasoning procedure: (1) collect evidence from multiple sources, (2) assess whether the evidence is sufficient or consistent to justify a decision, (3) resolve conflicts through debate mechanism, and (4) produce a verdict with traceable rationale.
 Overview
-Figure 1: High-level overview of our proposed AIFo.
-Our proposed framework, Agent-based Image Forensics (AIFo), is a training-free, LLM-based multi-agent system for AI-generated image detection, designed to mimic the procedural reasoning workflow of human forensic experts.
-As illustrated in Figure 1, the core of our framework [^1] We introduce the optional memory module in Section 5.
-consists of a Toolbox $\mathcal{T}=\{T_{1},T_{2},\dots,T_{n}\}$ containing $n$ different forensic tools, an Evidence Gatherer Agent $A_{EG}$, a Reasoning Agent $A_{R}$, two Debate Agent $A_{D1}$ and $A_{D2}$ and a Judge Agent $A_{J}$.
-Given an input image of unknown authenticity from the set of all images, $I\in\mathcal{V}$, the Evidence Gatherer $A_{1}$ first employs tools from the Toolbox.
-Specifically, the Toolbox includes four main types of forensic tools: (1) Image Reverse Search to locate visually similar images and retrieve their provenance information from online sources; (2) Metadata Extraction to recover embedded EXIF or generated metadata from the image file; (3) VLM-Based Analysis to perform semantic-level reasoning and contextual understanding of the image leveraging VLMs; and (4) Pre-Trained Classifiers to apply pre-trained models for binary classification of the image.
-Then, the Evidence Gatherer Agent executes each tool $T_{i}\in\mathcal{T}$ to produce a piece of evidence:
-$e_{i}=T_{i}(I),\quad T_{i}:\mathcal{V}\rightarrow\mathcal{E}_{i},$
-where $\mathcal{E}_{i}$ denotes the evidence space for tool $T_{i}$.
-For example, the metadata extraction tool may return detailed EXIF information embedded in the image file, such as "EXIF:Model": "Canon EOS 5D Mark IV" or "EXIF:LensModel": "EF24-70mm f/2.8L II USM", which can serve as important evidence for authenticity assessment.
-The complete evidence set is:
-$E=\{e_{i}\mid T_{i}\in\mathcal{T}\}.$
-Next, the Reasoning Agent $A_{R}$ examines the entire evidence set $\mathcal{E}$ and determines whether the evidence is sufficient and consistent enough to make a reliable decision.
-Evidence sufficiency refers to whether the collected evidence covers enough aspects of the input image, e.g., without missing key metadata or reverse search results.
-Evidence consistency measures the extent to which different pieces of evidence point toward the same conclusion, e.g., when the majority of evidence supports that the image is AI-generated.
-If the evidence is considered sufficient and consistent enough, the Reasoning Agent generates a final judgment $D\in\mathcal{D}$ and a human-readable explanation $R\in\mathcal{R}$:
-$A_{R}:\mathcal{E}\xrightarrow{}\mathcal{D}\times\mathcal{R}.$
-Otherwise, the system initiates a multi-round debate process.
-Two Debate Agents $A_{D1}$ and $A_{D2}$ take opposing stances (pro vs.
-contra) regarding whether the image is AI-generated or real.
-They exchange arguments over $n$ rounds, and a Judge Agent $A_{J}$ observes the debate history $H$ together with the tool-derived evidence $\mathcal{E}$, and produces the final judgment $D\in\mathcal{D}$ along with an explanation $R\in\mathcal{R}$:
-$A_{J}:(H,\mathcal{E})\rightarrow\mathcal{D}\times\mathcal{R}.$
-This hierarchical reasoning-debate framework ensures that decisions are made either directly from sufficient evidence or through a careful debating process that simulates human forensic reasoning when the evidence is inconclusive or conflicting.
-LLM-Based Agents
-The LLM-based agents act as the central reasoning and orchestration units of our multi-agent framework, responsible for applying forensic tools, evaluating the reliability of collected evidence, and making decisions based on existing evidence.
-All agents are instantiated from a LLM (GPT-4.1-2025-04-14), with specialized prompts defining their respective roles and responsibilities.
-The detailed prompts used for each agent are provided in Appendix A.
-The Evidence Gatherer has full access to the Toolbox and is responsible for invoking the forensic tools to collect evidence.
-Its decision-making process is guided by a carefully engineered prompt that contains three critical components: (1) a clear definition of the agent’s task, which is acting as an AI image forensics expert to determine whether the input image is AI-generated or real; (2) rigorous definitions of “AI-generated image” and “real image” (see Section 2.1) to ensure consistent interpretation during analysis; (3) a comprehensive description of all available tools, including their capabilities.
-With these explicit task definitions and resource descriptions, the LLM gains both the contextual understanding and the operational awareness to invoke the tools and collect results.
-The Reasoning Agent is responsible for conducting an initial assessment of the evidence collected by the evidence gatherer.
-It first evaluates whether the evidence set is sufficient and consistent enough.
-This evidence set check is implemented via a dedicated prompt, which instructs the LLM to internally assess the completeness (sufficiency) and consistency of the evidence and return a binary decision (true / false).
-If true, the reasoning agent proceeds to generate a final judgment and an explanation with another prompt that: (1) defines its responsibility to synthesize the collected evidence and produce a final binary judgment (AI-generated or real) with an accompanying explanation; (2) instructs the agent to evaluate the reliability of each evidence source, taking into account the confidence level, potential biases, and the trustworthiness of the generating tool; (3) emphasizes the importance of generating a human-readable, logically structured reasoning chain.
-If the evidence is insufficient or conflicting, the reasoning agent abstains from producing a final decision and instead triggers the multi-agent debate mechanism.
-The multi-agent debate mechanism is designed to resolve ambiguity and enhance the robustness of the final judgment when the evidence is inconclusive or conflicting.
-The Debate Agents are responsible for engaging in a structured multi-round debate when the evidence is deemed insufficient or inconsistent.
-The debate process lasts up to $n$ rounds.
-In the first round, the two debate agents are guided by designed prompts to take opposing stances: one agent provides arguments supporting that the image is AI-generated based on collected evidence, while the other provides arguments supporting that the image is real.
-In subsequent rounds, both agents are instructed to refine or strengthen their reasoning based on the arguments made by the opposing side in the previous round.
-Through this iterative exchange, the debate agents progressively sharpen their analyses and solve potential conflicts in the evidence interpretation.
-A Judge Agent is responsible for supervising the debate process.
-At the end of each round, the Judge Agent internally evaluates whether the debate has reached sufficient coverage and clarity to make a final decision.
-It may decide to terminate the debate early if the arguments are deemed sufficient, or allow the debate to proceed for additional rounds if further clarification is needed.
-Once the debate concludes, the Judge Agent generates the final judgment and an explanation.
-Unlike the Reasoning Agent, which relies solely on the collected evidence, the Judge Agent is guided to synthesize both the tool-derived evidence and the debate history.
-This design ensures that the final decision integrates multiple sources of evidence with structured deliberation, leading to a transparent and defensible forensic conclusion.
-Forensic Tools
-The forensic Toolbox is a collection of specialized modules that the Evidence Gatherer can invoke to analyze the input image. The tools can be broadly categorized into the following four classes:
-  • Reverse Image Search Tools: Assess image provenance by querying external online sources.
-  • Metadata Extraction Tool: Parse metadata information from the image file.
-  • Pre-Trained Classifiers Tool: Apply static models to assess authenticity.
-  • VLM-Based Reasoning Tool: Perform semantic-level analysis using VLMs.
-First, reverse image search tools are designed to identify the provenance and distribution history of input images by searching for exact matches or visually similar content across the Internet.
-These tools are particularly useful for determining image provenance, e.g., AI-generated images may appear on generative art platforms, while real images are more likely to be found on reputable news or photographic sites.
-We implement two complementary reverse image search tools.
-The first tool leverages the Google Cloud Vision API’s Web Detection service [7], which provides programmatic access to Google’s extensive image indexing capabilities.
-It primarily returns exact matches of the query image by analyzing its visual features and comparing them against Google’s web-scale image database.
-The API outputs structured data such as webpage titles, URLs, and contextual information where identical or highly similar images are found.
-However, the returned information can sometimes be sparse or insufficient for reliable judgment.
-To address this limitation and broaden the search scope, we implement a second approach inspired by Xu et al.
-[55] that simulates authentic human user behavior through web interface automation.
-This tool directly interacts with the Google Images search interface using Playwright, [^2] https://playwright.dev/ a modern web automation framework, to perform searches beyond the API’s restricted results.
-Unlike the first tool, this approach captures not only exact matches but also a wider range of visually similar images, thereby providing richer contextual evidence.
-It automatically uploads the input image, captures the results section from the returned webpage while filtering out advertisements and noise, and summarizes key information such as image match identification and the provenance of similar images.
-Metadata analysis tool focuses on extracting and analyzing technical metadata embedded within image files to identify authenticity markers that distinguish real photographs from AI-generated content.
-Digital cameras and imaging devices typically embed rich metadata (EXIF data) including camera settings, GPS coordinates, timestamps, and device-specific information that are often absent or inconsistent in synthetically generated images.
-For example, as shown in Table 1, a genuine photograph may contain entries like EXIF:Make and EXIF:Model indicating the camera manufacturer and model.
-They may also include realistic optical parameters or position data such as EXIF:LensInfo and Composite:GPSPosition that are indicative of a real-world capture event.
-In contrast, AI-generated images frequently lack such detailed metadata or contain synthetic traces.
-For instance, some AI images explicitly embed generation information or even the original text prompt used for synthesis, e.g., EXIF UserComment.
-These anomalies serve as strong indicators of AI generation.
-Our metadata extraction tool employs ExifTool, [^3] https://exiftool.org/ a comprehensive metadata manipulation library, to perform deep analysis of image files.
-Rather than extracting all available metadata fields, which can be overwhelming and include irrelevant information, our system implements a selective extraction strategy based on two key filtering mechanisms.
-First, we maintain a curated list of exact-match key fields $\mathrm{KEY\_FIELD\_EXACT}$ that have been empirically determined to provide strong authenticity signals.
-These include critical camera parameters such as focal length, aperture settings, ISO values, and camera manufacturer information.
-Second, we implement prefix-based filtering $\mathrm{KEY\_FIELD\_PREFIXES}$ to capture related metadata families.
-For example, all fields for camera manufacturer custom information are captured through the “MakerNotes” prefix.
-This dual-filtering approach ensures comprehensive coverage of relevant metadata while filtering out noise.
-The complete list of all key fields and prefixes is provided in the Appendix B.
-Table 1: Comparison of metadata signals for distinguishing real and AI-generated images.
-Metadata Field Example Value
-Real Image Signal
-EXIF:Make Canon
-EXIF:LensInfo 4.1 123 3.5 6.4
-Composite:GPSPosition 28.35 N, 81.59 W
-AI Image Signal
-JUMBF:Description AI Generated Image
-EXIF:UserComment "A full body of a cat..."
-Pre-trained classifiers tool leverages pre-trained models specifically fine-tuned for AI-generated image classification.
-These models have been trained on diverse datasets containing both authentic and AI-generated images, enabling them to learn discriminative features that distinguish between the two categories.
-We select the top five most downloaded classification models for AI-generated image detection available on Hugging Face [1]:
-  • haywoodsloan/ai-image-detector-deploy [3]
-  • Organika/sdxl-detector [12]
-  • legekka/AI-Anime-Image-Detector-ViT [4]
-  • Smogy/SMOGY-Ai-images-detector [13]
-  • NYUAD-ComNets/NYUAD_AI-generated_images
-    _detector
-     [11]
-Our implementation employs an ensemble approach that integrates multiple transformer models to enhance detection robustness and accuracy.
-The ensemble includes models trained on different synthetic image generation techniques, ensuring comprehensive coverage of various AI generation paradigms including GANs, diffusion models, and other neural synthesis methods.
-Each transformer model performs independent classification, producing probability distributions over the binary classification space.
-The final prediction score is obtained through a weighted voting mechanism, where each model’s contribution is scaled by its weight parameter.
-This weighted ensemble score is computed as:
-$\text{Prediction Score}=\frac{\sum_{i=1}^{N}w_{i}\cdot s_{i}}{\sum_{i=1}^{N}w_{i}},$
+Given an input image $I\in\mathcal{V}$, the objective is to produce a binary verdict $D\in\mathcal{D}=\{\texttt{AI},\texttt{Real}\}$ together with a rationale $R\in\mathcal{R}$.
+Rather than predicting $D$ directly from pixel-level features, the framework reasons over a set of heterogeneous evidence items $E=\{e_{1},e_{2},\dots,e_{m}\}$, where each $e_{k}$ is an evidence item produced from a distinct evidence channel (e.g., provenance, metadata, model predictions, semantic analysis).
+Each evidence item is represented as a structured record $e_{k}=\langle t_{k},\;s_{k},\;\xi_{k}\rangle$, where $t_{k}$ denotes the evidence type, $s_{k}\in\{\texttt{AI},\texttt{Real},\texttt{Unknown}\}$ indicates the qualitative indication suggested by the evidence source, and $\xi_{k}$ contains the structured output produced by the corresponding tool.
+In AIFo, the process of collecting evidence is carried out by an Evidence Gatherer Agent, which invokes a set of forensic tools and normalizes their outputs into the unified evidence representation above.
+The complete evidence set $E$ then serves as the input to a reasoning module, whose role is to qualitatively assess whether the evidence provides adequate support for a reliable decision and to identify agreements or contradictions across evidence items.
+The assessment such as evidence sufficiency and consistency are not computed as predefined metrics, but are instead explicitly reasoned about by a Reasoning Agent based on the available evidence and its contextual descriptions.
+Evidence sufficiency refers to whether the currently available evidence provides adequate support for a confident decision, and evidence consistency captures the degree to which different evidence sources support compatible conclusions.
+If the reasoning module concludes that the evidence is insufficient or contains unresolved conflicts, the framework defers commitment and invokes a structured conflict resolution procedure.
+Inspired by previous work [DLTTM24, LHJWWWYST24, LFLWD25, LDZHGLI24, SGDBP24], the framework adopts a multi-agent debate (MAD) mechanism to explicitly examine competing arguments rather than aggregating evidence.
+The debate is organized around two opposing hypotheses that the input image is AI-generated or real, and two agents are instantiated with complementary roles, each tasked with constructing the argument in support of one hypothesis using the same set of collected evidence.
+They exchange arguments over up to $n$ rounds, refining their reasoning based on prior exchanges at each round, and the debate continues until the supervising Judge Agent determines that one hypothesis is more defensible or the maximum number of rounds is reached.
+At termination, the Judge Agent synthesizes the debate history $H$ together with the original evidence $E$ to produce a final judgment $D\in\mathcal{D}$ and a corresponding explanation $R\in\mathcal{R}$: $A_{J}:(H,E)\rightarrow\mathcal{D}\times\mathcal{R}$.
+We instantiate the above procedural reasoning framework in AIFo, a training-free, LLM-based multi-agent system that maps each stage of the decision process to a specialized agent role.
+Each agent operates under a clearly defined role with restricted responsibilities and are guided by structured prompts.
+Evidence Providers and Tool Interfaces
+AIFo uses a modular toolbox of evidence providers.
+The forensic toolbox is a collection of specialized modules that the Evidence Gatherer Agent can invoke to analyze the input image.
+The tools can be broadly categorized into the following four classes.
+Reverse image search.
+Reverse image search tools provide provenance and online distribution of input images by searching for exact or visually similar matches across the internet.
+We employ two complementary search tools: one based on the Google Cloud Vision API for high-precision matches, and one web-automation method inspired by Xu et al.
+[XML24] to capture a broader set of visually similar results.
+Such provenance information can be highly indicative of authenticity, e.g., AI-generated images often appear on generative art platforms, whereas real images are more likely found on news or photography websites.
+Metadata Extraction.
+This tool extracts and analyzes embedded EXIF metadata (e.g., camera parameters, timestamps, GPS) to identify authenticity markers that differentiate real photographs from AI-generated images.
+We employ ExifTool for selective EXIF metadata extraction, focusing on key camera and capture-related fields while filtering irrelevant noise.
+The full list of key fields is provided in Appendix B.
+Pre-trained classifiers.
+We integrate multiple publicly available transformer-based detectors as an ensemble (details in Appendix C).
+Each model independently predicts the probability of AI generation and outputs an AI-generation score $s_{i}$.
+A final prediction score is obtained via a weighted voting scheme with equal model weights:
+$\text{Prediction Score}=\frac{\sum_{i=1}^{N}w_{i}\cdot s_{i}}{\sum_{i=1}^{N}w_{i}},$ (1)
 where $w_{i}$ represents the weight of the $i$-th model $\theta_{i}$, $s_{i}$ is the AI confidence score from $\theta_{i}$, and $N$ is the total number of loaded models.
-Each model is assigned an equal weight $w_{i}=1$ to ensure balanced contribution and avoid over-reliance on any individual model during inference.
-VLM-based reasoning tools employs VLMs to perform sophisticated visual analysis.
-These models leverage their extensive training on diverse image-text pairs to identify subtle visual patterns, artifacts, and inconsistencies that may indicate AI generation.
-Our implementation utilizes GPT-4.1 [8] to conduct detailed image analysis.
-A crafted prompt guides the model to focus on specific visual characteristics that are indicative of AI generation versus authentic photography.
-The complete prompt is provided in Table 15 in the Appendix.
-When the model identifies an image as AI-generated, it is prompted to provide detailed evidence in the form of specific visual artifacts.
-These include unnatural textures or patterns that deviate from expected material properties, inconsistent lighting or shadow directions that violate physical principles, anatomical errors in human or animal subjects, unusual distortions in object boundaries, text rendering abnormalities, symmetry issues, and contextual inconsistencies in background elements.
-For images classified as real, the model is instructed to explain the supporting visual characteristics, such as realistic anatomical proportions and coherent environmental context.
-Unlike the other forensic tools, VLMs not only enable binary classification but also provide detailed justifications and fine-grained visual analyses that enhance the interpretability of the final decision.
-Additionally, the model provides a confidence assessment (high, medium, or low) based on the strength and clarity of the observed evidence.
-The deterministic configuration (temperature = 0, seed = 42) ensures consistent and reproducible analysis results across multiple runs.
+We use equal weights for simplicity and this ensemble design enhances robustness and reduces overfitting to specific generators during inference.
+VLM-based analysis.
+A VLM-based tool conducts semantic image analysis to detect visual and contextual cues indicative of AI synthesis, examining artifacts such as lighting inconsistencies, unnatural textures, anatomical anomalies, and implausible contextual relationships.
+For real images, it identifies coherent structural, lighting, and contextual patterns.
+Beyond binary classification, the tool produces interpretable textual explanations and confidence estimates, enhancing the transparency and reliability of the judgments.
 Evaluation
 Dataset Construction
-To comprehensively evaluate our agentic framework, we consider both controlled and real-world scenarios.
-We therefore construct our benchmark dataset to cover two distinct settings: in-the-lab and in-the-wild.
-The in-the-lab setting consists of images collected from well-curated, controlled datasets commonly used in prior research, while the in-the-wild setting comprises images sourced from diverse, unconstrained online platforms, reflecting the complexity and unpredictability of real-world data.
-Our curated dataset comprises a total of 6000 images, evenly distributed across the two settings, with each containing 1500 AI-generated images and 1500 real images.
-For the in-the-lab setting, real images are sampled from well-established, curated datasets frequently used in image generation and detection research.
-Specifically, we sample 500 images each from Flickr30k [40], ImageNet [20], and the DIV2K dataset [16].
-The AI-generated counterparts are obtained from GenImage [61] and FakeBench [33].
-We sample 100 images from each of the 8 generative models included in GenImage, and 70 images from each of the 10 models in FakeBench, resulting in a total of 1500 AI-generated images.
-[…]
-Table 2: Dataset construction statistics.
-Source Datasets / Platforms Type # Images
-In-the-lab Setting
-Flickr30k [40] Real Image 500
-ImageNet [20] Real Image 500
-DIV2K [16] Real Image 500
-GenImage [61] AI Image 800
-FakeBench [33] AI Image 700
-In-the-wild Setting
-Holopix50k [26] Real Image 500
-Flickr [6] Real Image 500
-Wikimedia Commons [14] Real Image 500
-Lexica [9] AI Image 500
-NightCafe [10] AI Image 500
-Civitai [5] AI Image 500
-Experimental Setup
-Implementation Details.
-To evaluate the performance of our AIFo framework, we conduct experiments on the constructed benchmark dataset, comparing our agentic approach against a range of baseline methods.
-Our framework is implemented using LangGraph, [^4] https://langchain-ai.github.io/langgraph/ which is specifically designed for building stateful, multi-agent applications with LLMs.
-We select GPT-4.1 (gpt-4.1-2025-04-14 version) as the backbone for our LLM agents, and we set the temperature to 0 and the seed to 42 to ensure deterministic behavior across runs.
-Evaluation Metrics.
-We formulate the evaluation as a standard binary classification task, where the goal is to assess the framework’s ability to distinguish between AI-generated and real images.
-Specifically, we define correctly identifying an AI-generated image as a True Positive (TP) and correctly identifying a real image as a True Negative (TN).
-Based on these definitions, we compute the following standard evaluation metrics: Accuracy, Precision, Recall, and F1-score.
-In addition to evaluating the framework’s overall performance on the entire dataset, we further analyze its behavior separately on the in-the-lab and in-the-wild subsets to assess its robustness across controlled and real-world scenarios.
-Baseline Methods.
-To comprehensively assess the effectiveness of our proposed multi-agent framework, we compare it against a diverse set of representative baseline methods spanning both conventional and vision-language modeling paradigms.
-For traditional classifier-based approaches, we include CNNSpot [53], DE-FAKE [46], and PatchCraft [59], with their details discussed in Section 7.1.
-For VLM baselines, we adopt state-of-the-art multimodal models including GPT-4o and GPT-4.1, which have demonstrated advanced visual reasoning capabilities.
-During evaluation, each model is provided with the input image along with the prompt Is this a fake or real image?, and instructed to return a structured binary classification result.
-Results
 Table 1: Performance comparison of different methods on our benchmark dataset comprising three evaluation subsets: Overall, In-the-Lab, and In-the-Wild.
@@
 Method Acc Prec Rec F1 Acc Prec Rec F1 Acc Prec Rec F1
-CNNSpot [53] 0.5277 0.9826 0.0563 0.1066 0.5553 0.9882 0.1120 0.2012 0.5000 0.5000 0.0007 0.0013
-PatchCraft [59] 0.6517 0.7423 0.4647 0.5715 0.8123 0.8704 0.7340 0.7964 0.4910 0.4780 0.1953 0.2773
-DE-FAKE [46] 0.7142 0.6820 0.8027 0.7374 0.6720 0.6673 0.6860 0.6765 0.7563 0.6933 0.9193 0.7905
-GPT-4.1 [8] 0.9416 0.9913 0.8910 0.9385 0.9332 0.9932 0.8723 0.9288 0.9500 0.9895 0.9097 0.9479
-GPT-4o [39] 0.9483 0.9920 0.9038 0.9458 0.9537 0.9938 0.9130 0.9517 0.9428 0.9900 0.8947 0.9399
-AIFo w/o debate (ours) 0.9635 0.9922 0.9343 0.9624 0.9730 0.9924 0.9533 0.9725 0.9540 0.9921 0.9153 0.9521
+CNNSpot [WWZOE20] 0.5277 0.9826 0.0563 0.1066 0.5553 0.9882 0.1120 0.2012 0.5000 0.5000 0.0007 0.0013
+PatchCraft [ZXLQZ24] 0.6517 0.7423 0.4647 0.5715 0.8123 0.8704 0.7340 0.7964 0.4910 0.4780 0.1953 0.2773
+DE-FAKE [SLYZ22] 0.7142 0.6820 0.8027 0.7374 0.6720 0.6673 0.6860 0.6765 0.7563 0.6933 0.9193 0.7905
+GPT-4o [O23] 0.9483 0.9920 0.9038 0.9458 0.9537 0.9938 0.9130 0.9517 0.9428 0.9900 0.8947 0.9399
 AIFo (ours) 0.9705 0.9920 0.9487 0.9698 0.9740 0.9917 0.9560 0.9735 0.9670 0.9923 0.9413 0.9661
-Table 3 presents the comprehensive performance comparison of our multi-agent framework against five baseline methods across three evaluation settings.
-A detailed breakdown of their performance is provided in Table 17 in the Appendix.
-These results demonstrate the superior effectiveness of our proposed approach in AI-generated image detection under both controlled and real-world conditions.
-Traditional Classifier Baselines.
-The traditional methods show varied performance patterns.
-CNNSpot [53] exhibits poor overall detection performance, close to random guessing.
-Its extremely low recall and F1-score indicate that it fails to effectively detect AI-generated images.
-DE-FAKE [46] achieves moderate performance, with an accuracy of 0.7142 and an F1 score of 0.7374, slightly outperforming PatchCraft [59].
-These results highlight the inherent limitations of traditional classification methods in coping with rapidly evolving and continuously updated diverse AI-generated content in complex real-world scenarios.
-Vision-Language Model Baselines.
-Both GPT-4.1 and GPT-4o demonstrate strong performance, with accuracy exceeding 0.94 and F1-scores above 0.93.
-However, the recall rates for both models are relatively lower compared to their precision, indicating that they still miss a significant portion of AI-generated images.
-This suggests challenges remain for these models in detecting certain types of AI-generated content.
-Our agent framework addresses this limitation through the multi-agent architecture, which combines the strengths of multiple specialized tools to achieve more reliable detection.
-AIFo Framework Performance.
-Compared to all baseline methods, our AIFo framework consistently achieves the best performance across almost all metrics and evaluation settings.
-In the overall evaluation, AIFo attains an accuracy of 0.9705 and an F1-score of 0.9698, surpassing the strongest baseline GPT-4o by absolute margins of 2.22% in accuracy and 2.40% in F1-score.
-Moreover, this improvement is observed consistently in both in-the-lab and in-the-wild data, which span over 20 generative models, demonstrating the strong generalization capability of our framework.
-Through a training-free paradigm, our method is minimally affected by the evolution and iteration of generative models.
-Impact of Multi-Agent Debate Mechanism.
-We also evaluate the impact of disabling the multi-agent debate mechanism on the performance.
-Notably, even without the debate mechanism, our AIFo framework still significantly outperforms GPT-4o.
-With the debate mechanism incorporated, our framework’s accuracy is further improved, particularly in in-the-wild setting, where it surpasses GPT-4o by 2.4%.
-These results demonstrate that the debate mechanism serves as an effective refinement stage, resolving potential reasoning uncertainties caused by conflicting evidence and enhancing the reliability of the final decision.
-Laboratory vs.
-Wild Environment Analysis.
-A comparison between controlled (in-the-lab) and real-world (in-the-wild) environments reveals key differences.
-Traditional classifiers often fail to generalize: for example, PatchCraft’s F1-score drops from 0.7964 in the laboratory to 0.2773 in the wild, and CNNSpot nearly collapses entirely.
-This highlights the difficulty of transferring models trained on curated datasets to diverse real-world content.
-In contrast, our agent framework achieves consistently strong results, with F1-scores of 0.9735 in the laboratory and 0.9661 in the wild, across more than 20 generative models.
-These findings confirm the framework’s ability to generalize beyond controlled datasets and effectively handle the diversity of real-world AI-generated content.
+To comprehensively evaluate our multi-agent framework under both controlled and open-world conditions, we construct a benchmark dataset covering two complementary settings: in-the-lab and in-the-wild.
+The in-the-lab setting consists of well-curated datasets commonly used in prior work, where image sources and generation processes are relatively controlled.
+In contrast, the in-the-wild setting comprises images collected from unconstrained online platforms, reflecting the diversity, noise, and uncertainty encountered in real-world forensic scenarios.
+The benchmark contains 6,000 images in total, evenly split between the two settings, with 1,500 real and 1,500 AI-generated images per setting.
+For the in-the-lab setting, real images are sampled from Flickr30k [PWCCHL15], ImageNet [DDSLLF09], and DIV2K [AT17] (500 images each).
+AI-generated images are obtained from GenImage [ZCYHLLTHHW23] and FakeBench [LLWLWRL24], by sampling 100 images from each of the eight models in GenImage and 70 images from each of the ten models in FakeBench, yielding 1,500 AI-generated samples.
+For the in-the-wild subset, real images are collected from publicly available online sources, including Flickr, Wikimedia Commons, and the Holopix50k dataset [HKURGOL20].
+To ensure diversity, we select ten keywords: animal, building, food, indoor, landscape, person, plant, snow, sport, transportation and water.
+Images from Flickr and Wikimedia Commons are retrieved via keyword search and randomly sampled from the results.
+For Holopix50k, which lacks explicit semantic labels, we use BLIP [LLXH22] to categorize images into the same keyword set before random sampling.
+For each keyword and each source, we sample 50 images, yielding a balanced collection of real-world photographs.
+The corresponding AI-generated images are sourced from three popular generative art platforms: Lexica, NightCafe, and Civitai.
+Images from these platforms are generated by a wide range of text-to-image models, including DALL·E [RDNCC22], Stable Diffusion [RBLEO22], SDXL [PELBDMPR23], StyleGAN [KLAHLA20], and numerous community fine-tuned variants.
+We use the same set of ten keywords to retrieve AI-generated images and randomly sample 50 images per keyword from each platform.
+As summarized in Appendix D, the resulting benchmark spans over 20 generative models, enabling evaluation under both controlled and open-world conditions.
+Experimental Setup
+Implementation Details..
+We evaluate AIFo under a unified decision protocol that reflects its full procedural reasoning pipeline.
+The framework is implemented using LangGraph, which supports stateful multi-agent workflows.
+All agents share the same LLM backbone (e.g., GPT-4o), ensuring consistent reasoning capacity across different agent roles.
+Unless otherwise specified, all agents are instantiated with temperature set to 0 and a fixed random seed (42) to reduce stochastic variation in reasoning traces.
+All external tool outputs, including reverse image search results and metadata extraction, were cached at evaluation time and treated as fixed inputs during benchmarking to ensure reproducibility.
+Evaluation Metrics..
+We evaluate performance using standard binary classification metrics, including Accuracy, Precision, Recall, and F1-score, where AI-generated images are treated as positive samples.
+While these metrics provide a high-level summary of detection effectiveness, they do not capture how or why a decision is reached.
+We therefore include targeted analyses that examine intermediate decision dynamics, such as when additional reasoning is triggered and how often the final verdict differs from the initial assessment.
+Baseline Methods..
+We compare AIFo against representative baseline approaches spanning both conventional classifiers and VLMs.
+For classifier-based methods, we include CNNSpot [WWZOE20], DE-FAKE [SLYZ22], and PatchCraft [ZXLQZ24].
+For VLM baselines, we adopt the same LLM backbone as used by the agents (e.g., GPT-4o) to ensure a fair comparison.
+Each VLM receives the input image together with a binary classification prompt.
+We experiment with multiple prompt techniques for the VLM baselines such as Chain-of-Thought (CoT) [WWSBIXCLZ22] and report the best results, to provide a competitive reference for evaluating the effectiveness of our framework.
+Results
+Overall Performance.
+Table 1 reports the overall performance of AIFo and baseline approaches across three evaluation settings.
+The results in this table are obtained using GPT-4o as the shared LLM backbone for all agents.
+These metrics summarize the outcome of the complete decision process, in which a verdict may be produced either directly from initial evidence assessment or after invoking conflict-aware procedural reasoning.
+A detailed per-setting breakdown is provided in Appendix E.
+Traditional classifiers (CNNSpot [WWZOE20], DE-FAKE [SLYZ22], PatchCraft [ZXLQZ24]) perform poorly, with limited ability to generalize beyond curated datasets.
+GPT-4o achieves strong overall accuracy but tends to rely predominantly on visual semantics, leading to lower recall in cases where AI-generated images closely mimic real photographic patterns.
+Across all evaluation settings, AIFo achieves competitive or superior overall performance compared to all baselines.
+In the full evaluation, it attains an accuracy of 0.9705 and an F1-score of 0.9698, exceeding GPT-4o by 2.22% in accuracy and 2.40% in F1-score.
+Figure 2: Examples of our agent framework’s decision-making process.
+Table 2: Ablation Study of AIFo Components with GPT-4o Backbone
+Method Acc Prec Rec F1
+GPT-4o (direct) 0.9483 0.9920 0.9038 0.9458
+GPT-4o + CoT 0.9510 0.9906 0.9107 0.9489
+AIFo w/o Tools 0.9525 0.9913 0.9130 0.9505
+AIFo w/o MAD 0.9635 0.9922 0.9343 0.9624
+AIFo 0.9705 0.9920 0.9487 0.9698
 Quantitative and Qualitative Analysis.
-To better understand where AIFo achieves improvements over GPT-4o, we analyze all samples misclassified by GPT-4o but correctly identified by AIFo.
-We find that AIFo correctly identifies 136 more AI-generated images compared to GPT-4o out of the total 6,000 samples.
-Among these corrected cases, approximately 38% are attributed to decisive reverse image search evidence, primarily from in-the-wild sources.
-Another 35% are resolved through metadata signals, which prove particularly effective for images containing detailed camera parameters or generation prompts embedded by community diffusion models.
-The remaining 27% rely on transformer-based classifier results as supporting evidence.
-In approximately 30% of all corrected cases, AIFo invokes the debate mechanism to resolve ambiguous or conflicting evidence, ultimately overturning the initial VLM-based assessment.
-To further demonstrate the effectiveness of AIFo, we present qualitative examples in Figure 2 that show how the system integrates multiple sources of evidence and, when necessary, employs a debate mechanism to reach accurate conclusions.
-In the first real-image case, the Evidence Gatherer agent successfully collects evidence from four forensic tools, including image sources, camera parameters, VLM analysis, and classifier scores.
-The Reasoning Agent judges the evidence sufficiency and, by synthesizing all sources, correctly concludes that the image is real.
-In the second AI-generated case, the realistic appearance misleads the baseline GPT-4o and the VLM tool, but other tools indicated an AI origin, creating conflicting signals.
-In this case, the Reasoning Agent judges that the evidence is inconsistent for a direct decision and instead triggers the debate mechanism.
-The pro-debate agent and con-debate agent present arguments supporting their respective stances, and the decisive argument came from the pro side, which highlighted the exact match of the image found on an AI platform.
-Ultimately, the Judge Agent weighs the conflicting evidence and reaches the correct conclusion that the image is AI-generated.
-These cases highlight two key strengths of our framework: its ability to override misleading visual evidence using stronger provenance signals, and its capacity to reconcile conflicting evidence through structured debate.
-These capabilities are essential for reliable detection in both benchmarks and real-world scenarios.
-Figure 2: Examples of our agent framework’s decision-making process, demonstrating diverse evidence integration across different image types and sources.
+To understand where the performance gains of AIFo originate, we analyze test samples misclassified by GPT-4o but correctly classified by AIFo.
+Among 6,000 samples, AIFo correctly identifies over 140 additional AI-generated images.
+Approximately 64% of these corrections are attributed to external evidence unavailable to single VLM inference.
+Roughly 34% of the corrected samples involve insufficient or conflicting signals across different evidence channels, which trigger the multi-agent debate process.
+In such cases, initial assessments are revised through structured evaluation of competing hypotheses, yielding a corrected final verdict by the Judge Agent.
+Table 2 further shows that removing either external evidence tools or the multi-agent debate mechanism consistently degrades performance compared to the full framework, confirming that both components contribute meaningfully to the overall gains.
+Qualitative examples in Figure 2 further illustrate this process.
+In the first real-image case, multiple evidence sources provide consistent support for authenticity, allowing the framework to reach a direct verdict without invoking further conflict resolution stage.
+In the second case, the realistic appearance misleads the baseline GPT-4o and the VLM tool.
+However, conflicting provenance and classifier evidence induce the framework to defer commitment and invoke multi-agent debate.
+Through structured argumentation, the system identifies the inconsistency in the assessment and ultimately produces the correct AI-generated verdict.
+Overall, this analysis shows that the gains of AIFo come from two complementary factors: the ability to override misleading visual cues using stronger non-visual evidence, and the capacity to revise initial judgments through explicit reasoning when evidence is insufficient or contradictory.
 Tool Reliability and Decision Pattern Analysis.
-To gain deeper insights into the internal decision-making processes of our multi-agent framework, we conduct a quantitative analysis of the reliability and usage patterns of individual tools, as illustrated in Figure 3.
-We define a consistent judgment as a case in which the agent’s final decision is in agreement with the evidence provided by a given tool.
-The reliability rate of a tool is then computed as the ratio of its consistent judgments to the total number of valid (i.e., informative) pieces of evidence it produces.
-A higher reliability rate indicates that the agent frequently trusts and aligns with that tool’s output, suggesting its potential influence on the overall decision-making process.
-We also compute the coverage rate of each tool, defined as the proportion of cases in which the tool provides valid evidence relative to the total number of agent decisions.
-From 3(a) and 3(b), metadata extraction and the first reverse search tool focused on exact matches achieve the highest reliability rates, meaning their evidence plays a significant role in shaping the agent’s final decisions when available.
-However, their lower coverage rates show they are only relevant to a subset of images.
-This is because in our benchmark dataset, only a small fraction of images contain rich EXIF metadata or have provenance information retrievable using the first reverse search tool.
-These tools thus act as high-precision, low-frequency decision anchors, exerting strong influence when available but not contributing in all cases.
-In contrast, the VLM analysis tool, pre-trained classifiers, and the second reverse search tool specialized in similar image retrieval achieve nearly full coverage.
-While the VLM’s visual reasoning remains relatively reliable, the outputs of the second reverse search tool and the classifiers are less dependable.
-This suggests that the agent will internally assign lower weights to these pieces of evidence.
-Hence, these tools serve as broad-coverage, moderate-confidence sources of evidence, providing additional context and support for the final decision but not dominating the reasoning process.
-Overall, the agent appears to follow a tiered weighting strategy: prioritizing high-reliability tools when available, while leveraging high-coverage tools to maintain decision robustness.
-(a) Tool reliability rates
-(b) Tool coverage rates
-Figure 3: Analysis of individual tool contributions to the agent framework: (a) reliability rates measuring agent trust in each tool’s evidence, and (b) coverage rates showing the proportion of decisions where each tool provides informative evidence.
+To better understand the internal decision-making processes of our framework, we analyze the reliability and coverage of individual forensic tools.
+We define reliability as the proportion of cases where a tool’s evidence aligns with the agent’s final decision, and coverage as the proportion of decisions for which a tool provides valid evidence.
+As shown in Figure 3, metadata extraction and the first reverse search tool achieve the highest reliability, indicating that their evidence strongly influences final decisions when available.
+However, their relatively low coverage indicates that such decisive evidence is only present in a subset of cases.
+These tools therefore function as high precision but sparse evidence sources.
+In contrast, the VLM-based analysis, pre-trained classifiers, and the second reverse search tool provide near-complete coverage across the dataset.
+Among them, VLM-based analysis contributes the most essential visual evidence and serves as the primary source of information in the majority of cases.
+This observation is further supported by a leave-one-out analysis shown in Figure 4, where removing the VLM tool leads to the most pronounced performance degradation, with accuracy dropping below 0.85 and recall below 0.70.
+By comparison, disabling metadata extraction, reverse image search, or pre-trained classifiers results in only modest performance degradation, indicating that these tools provide complementary rather than dominant signals.
+Instead of relying on any single tool, AIFo leverages procedural reasoning to balance these heterogeneous signals.
+Highly reliable but low coverage evidence is emphasized when present, while high coverage but noisier evidence provides contextual support when stronger signals are unavailable.
+This behavior emerges from the reasoning process itself and is not explicitly hard-coded into the system.
 Inference Efficiency and Cost Analysis.
-While AIFo achieves high detection accuracy, its architecture introduces additional computational cost compared to single model such as GPT-4o.
-To assess the efficiency of our framework, we measure the end-to-end latency and token usage for processing a single image.
-Table 4 summarizes the inference latency and token consumption per image for different methods.
-On average, AIFo requires 40.08 seconds per image, approximately 7.5 times slower than GPT-4o’s 5.31 seconds.
-The token usage also increases significantly, with AIFo consuming an average of 5230.86 tokens per image, compared to GPT-4o’s 715.05 tokens.
-The majority of the additional latency and token usage originates from the multi-round llm invocation and the execution of external tools.
-Despite the increased computational cost, AIFo achieves a consistent 2-3 % improvement in accuracy and over 4 % gain in recall.
-Meanwhile, it is important to note that AIFo operates in a training-free manner, eliminating the need for costly model retraining or fine-tuning as generative models evolve.
-Most importantly, unlike single model reasoning that relies solely on visual analysis, our framework integrates diverse sources of evidence, offering more verifiable interpretability that aligns with human forensic reasoning.
+The procedural reasoning framework incurs higher computational cost than VLM baselines due to multi-stage reasoning, multi-agent interaction, and external tool invocation.
+As summarized in Table 3, AIFo requires an average of 40.08 s and 5.2k tokens per image, corresponding to approximately $7.5\times$ and $7.3\times$ the cost of GPT-4o, respectively.
+While this overhead is non-negligible, it should be interpreted in the context of the framework’s adaptive decision process.
+As shown in earlier analyses, only a subset of samples require conflict resolution reasoning, while the majority are resolved after the initial evidence assessment.
+More importantly, the additional computation enables capabilities that single-shot visual reasoning cannot provide.
+By integrating provenance, metadata, and model-based evidence, AIFo offers verifiable decision rationales and improved reliability in ambiguous cases, which are critical in high-stakes applications.
+These results highlight a fundamental trade-off between efficiency and decision robustness, and suggest that our procedural reasoning is most appropriate when interpretability and reliability are prioritized over real-time constraints.
+Figure 3: Analysis of individual tool contributions to the agent framework.
+Figure 4: Performance degradation when each tool is disabled from the framework.
 Table 3: Average inference latency and token usage per image.
@@
 GPT-4o 5.31s 715.05
-AIFo (w/o debate) 25.43s 2728.29
+AIFo (w/o MAD) 25.43s 2728.29
 AIFo 40.08s 5230.86
-Robustness Analysis
-To assess the robustness of our multi-agent detection framework under realistic conditions, we evaluate its performance on perturbed versions of the dataset.
-The goal of this analysis is to determine whether the framework can maintain consistent accuracy when the input images undergo quality degradation.
-We apply a series of standard image degradation techniques, including: Blurring.
-We apply Gaussian blur with a fixed radius of 2, simulating defocus or motion blur; Sharpening.
-We enhance image sharpness using a sharpening factor of 2.0, which may alter local edge distributions; Gaussian Noise.
-We add Gaussian noise with mean 0 and variance 2 to simulate sensor-level or environmental noise.
-These image transformations represent common distortions encountered in real-world image acquisition or compression pipelines.
-For each perturbation type, we measure the detection performance on the perturbed images and compare the results against the best baseline, GPT-4o.
-Table 5 presents the comprehensive performance comparison between our AIFo framework and the GPT-4o baseline across three types of image perturbations: Blurring (Blu), Gaussian noise (Noi), and sharpening (Sha).
-The results demonstrate that our framework maintains superior performance under these robustness testing conditions.
-Table 5: Performance comparison of AIFo vs GPT-4o under different image perturbations on the overall dataset. AIFo superior results are highlighted in bold.
+Table 4: AIFo vs. GPT-4o performance under image perturbations. AIFo ’s superior results are in bold.
 GPT-4o AIFo (Ours)
@@
 Sha 0.9410 0.9926 0.8887 0.9377 0.9670 0.9902 0.9433 0.9662
-Our AIFo framework consistently outperforms GPT-4o across all perturbation types and most evaluation metrics.
-Under blurring conditions, both GPT-4o and AIFo experience noticeable performance degradation, as visual analysis tools are inherently sensitive to loss of a great amount of image pixel features.
-Nevertheless, AIFo still achieves an accuracy of 0.9047, outperforming GPT-4o’s 0.8818.
-In contrast, under noisy and sharpened conditions, AIFo demonstrates remarkable robustness: its accuracy remains at 0.9710 and 0.9670 respectively, with only minimal decline compared to the clean setting, and consistently surpasses GPT-4o.
-This robustness can be attributed to the fact that, apart from pre-trained classifiers, the majority of tools integrated in our framework are largely insensitive to the perturbations, thereby stabilizing overall performance.
-Evasive Attack Analysis
+Robustness Analysis
+To evaluate robustness of our framework under realistic degradations, we test its performance on perturbed versions of the dataset.
+Three common distortions are applied: Gaussian blur (radius = 2) to simulate defocus, Gaussian noise (mean = 0, variance = 2) to mimic sensor noise, and sharpening (factor = 2.0) to simulate edge distributions.
+These perturbations reflect typical degradations in real-world imaging and compression pipelines and we compare the performance against the best baseline, GPT-4o.
+As shown in Table 4, both AIFo and GPT-4o experience performance degradation under blur due to the loss of fine-grained visual details.
+However, AIFo consistently maintains higher accuracy across all perturbation types.
+We observe that perturbations increase disagreement between visual evidence and other evidence sources, thereby triggering debate mechanism more frequently.
+This allows the framework to rely less on degraded visual cues and instead emphasize complementary non-visual evidence when available.
+Under noise and sharpening, where visual distortions are less severe, AIFo exhibits only minor performance drops relative to the clean setting.
+This behavior suggests that procedural reasoning provides a mechanism for adapting to changes in evidence reliability.
+Overall, these results indicate that robustness in AIFo emerges from its ability to mediate heterogeneous evidence under varying degrees of visual uncertainty.
 To evaluate the resilience of AIFo against evasive attacks, we simulate two representative attack scenarios that target the framework’s key evidence sources.
@@
 None (clean) 0.9705 0.9920 0.9487 0.9698
-Reverse-search manipulation 0.8971 0.8690 0.9353 0.9010
+Reverse search manipulation 0.8971 0.8690 0.9353 0.9010
 Metadata forgery 0.8702 0.8399 0.9147 0.8757
-Ablation Study
-To investigate the individual contributions of different tool categories within our multi-agent framework, we conduct ablation experiments to quantify the impact of each of the four tool types on overall detection performance.
-To assess the importance of each category, we adopt a leave-one-out strategy: in each ablation run, we disable one tool category while keeping the others unchanged, and evaluate the resulting performance on the full benchmark dataset.
-Meanwhile, we temporarily disable the debate mechanism to better isolate and focus on the contributions of the tools themselves.
-By observing the performance drop associated with the removal of each tool category, we can quantify its relative contribution to the framework’s overall effectiveness.
-This analysis provides insight into which components are most essential for robust and reliable detection, and helps guide future efforts in optimizing or simplifying the system without substantial loss in performance.
-Figure 4 presents the performance degradation observed when each tool category is disabled.
-The results clearly demonstrate that the vision-language model (VLM) tool contributes the most critical evidence to the overall detection pipeline.
-Removing the VLM tool causes the largest performance degradation across all metrics, with accuracy dropping below 0.85, and recall falling below 0.70.
-This indicates that direct image analysis from state-of-the-art VLMs remains the main source of evidence for the reasoning agent.
-Disabling any of the metadata extraction, reverse image search, or pre-trained classifier modules leads to a slight decrease in both accuracy and F1 score.
-This indicates that the evidence provided by these tools offers complementary information that enhances the capability of the reasoning agent.
-Notably, even without these modules, our framework still surpasses the GPT-4o baseline, confirming that these tools contribute complementary signals that strengthen the robustness of the overall reasoning process.
-Overall, the ablation study validates the necessity of a diverse Toolbox for reliable detection.
-These findings highlight that performance gains in our multi-agent framework arise not from a single dominant component, but from the effective integration of cross-source and complementary evidence sources.
-Figure 4: Performance degradation when each tool is disabled from the framework.
 Takeaway
-Our evaluation demonstrates the advantages of the proposed multi-agent framework: it consistently achieves the highest detection performance on five benchmark datasets in the lab and in-the-wild images from six online platforms, covering AI-generated content produced by at least 20 models.
-The framework also shows stronger robustness than the best baseline, GPT-4o, against three common perturbations.
-We find that our framework can dynamically weigh the confidence of evidence gathered from each tool: when metadata and reverse search tools provide valid evidence, it tends to prioritize them; otherwise, it falls back on VLM analysis and traditional classifiers.
-The debate module further addresses conflicting evidence and prevents the framework from over-relying on a single source of evidence.
-Altogether, the framework demonstrates greater effectiveness and robustness compared to single-model approaches for AI-generated image detection.
+This evaluation demonstrates that procedural reasoning offers distinct advantages for AI-generated image detection, particularly in scenarios where evidence is incomplete, ambiguous, or conflicting.
+By explicitly assessing evidence sufficiency and resolving conflicts through structured reasoning, the framework is able to revise misleading initial judgments and produce more stable and interpretable decisions.
+At the same time, these benefits come with increased computational cost, underscoring a fundamental trade-off between efficiency and decision robustness.
+Our results suggest that procedural, agent-based reasoning is most appropriate for forensic and high-stakes applications, where reliability, and auditability are prioritized over real-time inference.
+Together, these findings support that AI-generated image detection is better viewed as an evidence-driven decision process rather than a purely predictive task, and highlight procedural reasoning as a promising direction for addressing the limitations of existing detectors.
 Memory-Augmented Reasoning
 Motivation and Design
-Figure 5: Overview of the memory-augmented reasoning module.
-Our current designed detectors operate as stateless processes, analyzing each image independently without leveraging historical context or learning from past decisions.
-However, forensic analysis is inherently iterative and cumulative, where insights from previous cases can inform future judgments [25].
-To address this limitation, we introduce a memory-augmented reasoning module that enables the framework to maintain a knowledge base of past cases and decisions.
-As illustrated in Figure 5, the module captures, indexes, and retrieves relevant historical context to inform current detection decisions.
-The designed knowledge base systematically collects both successful and failed detection cases, storing image embedding vectors together with their associated classification results, collected evidence, and reasoning processes.
-For failed cases, it additionally stores a reflective analysis that captures the causes of misclassification.
-When analyzing a new image, the system performs similarity-based retrieval over this knowledge base to identify relevant historical cases.
-We hypothesize that the introduction of similar historical cases enables the agent to learn from experience and potentially adjust the weights of evidence from different forensic tools based on past performance patterns.
-Therefore, the enhanced system first queries its accumulated experience to identify potentially relevant historical cases rather than approaching each image as an isolated classification problem.
-Implementation Details
-Knowledge Base Dataset.
-To avoid data leakage, we construct the knowledge base image repository using a completely separate set of images from those used in the main benchmark dataset described in  Section 4.1.
-Following the same collection and sampling methodology as the benchmark dataset, we curate 600 new images for the knowledge base, containing 300 AI images and 300 real images.
-Building Knowledge Base.
-Our knowledge base system is built on a hybrid architecture that combines vector similarity search for content retrieval with structured metadata storage for case management and analysis.
-The implementation leverages CLIP [42] embeddings to capture high-level semantic and visual features that enable effective similarity matching across images.
-Each image processed by the framework undergoes feature extraction using the CLIP-ViT-B/32 model, generating 512-dimensional dense vector representations.
-These embeddings serve as the primary indexing mechanism, enabling efficient similarity-based retrieval of historically relevant cases.
-The knowledge base maintains separate indices for successful and failed detection cases.
-For successful cases, the system records the complete set of evidence used in decision-making and the full analytical process.
-For failed cases, the system records all the above information and additionally stores a structured reflection, providing deeper insight into the causes of misclassification.
-The reflection is generated by GPT-4.1 to examine the complete analytical pathway and identify potential failure points.
-This process produces insights covering evidence misinterpretation, tool reliability assessment, and reasoning inconsistencies.
-These generated reflections create a systematic learning mechanism that helps the system avoid repeating similar analytical errors.
-Finally, the knowledge base module is integrated seamlessly as one extra forensic tool.
-During the evidence collection phase, the module is activated.
-Inference.
-Upon receiving a new image, the system computes its CLIP embedding and performs a similarity search over the knowledge base indices to retrieve the top-$k$ most relevant historical cases, with defaults of $k=1$ and a similarity threshold of $0.85$ to ensure only sufficiently close matches are returned.
-The retrieved cases are then passed to the reasoning agent as additional context, allowing it to leverage past experiences to inform the current analysis.
+The core procedural reasoning framework operates in a stateless manner, analyzing each image independently without access to prior cases.
+While this design ensures fairness and prevents information leakage in benchmark evaluation, human forensic reasoning is often cumulative, drawing on past experiences to inform judgments in difficult or ambiguous cases [HRMP25].
+Motivated by this observation, we explore a memory-augmented reasoning module as an optional component that provides contextual reference to past cases.
+The memory module maintains a knowledge base of previously analyzed cases, including their visual embeddings, collected evidence, final decisions, and reasoning traces.
+For failure cases, it additionally stores structured reflections generated by LLM that describe potential causes of error, such as unreliable evidence sources or misinterpretation of visual cues.
+During inference, the memory module retrieves semantically similar cases using CLIP-based embedding similarity.
+These retrieved cases are passed to the reasoning agent as supplementary context, allowing it to reflect on similarities and differences between the current input and past cases.
+The agent may choose to incorporate this information when reassessing evidence reliability or resolving conflicts, but is not required to follow past decisions.
+This design reduces the risk of systematic bias or error propagation, as memory serves as a reference instead of a prescriptive rule.
+To avoid data leakage, the knowledge base is constructed from a separate set of 600 images (300 real, 300 AI-generated) distinct from the benchmark dataset.
+Each image is embedded using CLIP-ViT-B/32 [RKHRGASAMCKS21], and indexed for similarity retrieval.
+The memory module is integrated as an additional forensic tool within the framework and activated during evidence collection.
+When analyzing a new image, it retrieves the top-$k$ most relevant past cases ($k=1$ by default), which are passed to the reasoning agent as additional context.
 Results Analysis
-We did not incorporate the memory module in the main evaluation, as it requires ground-truth labels during the experience accumulation phase.
-Allowing such label-dependent learning at test time would introduce unfair advantages to our framework, compared to baselines that do not access these labels.
-Instead, to highlight the potential of the memory module, we directly focus on failure cases from the main evaluation.
-These are the scenarios where the knowledge base and additional experience are most needed.
-To more directly assess its effectiveness, we conduct a targeted evaluation focusing on failure cases from the original framework.
-Specifically, we identify 50 misclassified images from the benchmark set for which the knowledge base contained semantically similar counterparts.
-We then re-evaluated these images with the knowledge base module enabled, examining whether the additional contextual evidence could help correct the prior misjudgments.
-Table 7 summarizes the results by different image types and settings, showing the number of errors successfully corrected after incorporating knowledge base evidence.
-The results demonstrate that over 40% of error cases are successfully corrected when the memory module is employed.
-Furthermore, the example illustrated in Figure 6 highlights how retrieved similar cases influenced the reasoning process.
-When presented with similar past cases, the agents internally reweighted the credibility of different forensic tools, reinterpreted conflicting evidence, and ultimately arrived at the correct conclusion.
-These observations suggest that the memory-augmented reasoning offers substantial benefits in failure recovery and demonstrates strong potential for enhancing adaptive reasoning in forensic scenarios.
-Table 7: Number of errors before and after incorporating similar case history. FP: real images misclassified as AI images; FN: AI images misclassified as real images.
+We do not include the memory-augmented module in the main evaluation, as constructing the memory knowledge base requires access to ground-truth labels during the experience accumulation phase.
+Including such label dependent information at test time would introduce an unfair advantage compared to baseline methods.
+Instead, we evaluate the memory module through a controlled failure recovery analysis, which reflects its intended use as an additional mechanism for difficult cases.
+Specifically, we analyze 50 misclassified images from the main benchmark for which semantically similar cases exist in the knowledge base.
+These cases represent scenarios where the stateless reasoning fails and additional contextual reference may be beneficial.
+After enabling memory-augmented reasoning, over 40% of these previously misclassified samples are corrected (Table 6).
+Qualitative examples in Figure 5 illustrate how retrieved past cases support reflective reasoning.
+These observations suggest that the memory-augmented reasoning offers substantial benefits in failure recovery and strong potential for enhancing adaptive reasoning.
+Table 6: Number of errors before and after incorporating similar case history. FP: False Positive Cases; FN: False Negative Cases.
 In-the-Lab In-the-Wild
@@
 After 1 6 0 22 29
-Figure 6: Example of memory-augmented reasoning showing how historical cases influence the agent’s decision-making process and evidence weighting.
+Figure 5: Example of memory-augmented reasoning illustrating the impact of similar historical cases on decision-making.
 Discussion and Limitations
@@
 Adversarial Metadata Manipulation.
-Our framework lies in its reliance on image EXIF metadata as a key source of forensic evidence.
+The framework’s reliance on image EXIF metadata as a key source of forensic evidence is a limitation.
 However, adversaries could potentially manipulate image metadata to mislead the detection system.
@@
 Such metadata spoofing attacks could compromise the reliability of our metadata extraction tool, which currently shows high reliability rates in our evaluation.
-Related Work
-Fake Image Detection
-Early approaches to AI-generated image detection mainly rely on training image classifiers using machine learning, often using datasets produced by specific generative models.
-Over time, research has shifted towards more generalizable and robust detection strategies, including leveraging the visual capabilities of large multimodal models (LMMs) such as CLIP [42].
-CNNSpot [53] is one of the first works to propose a “universal” detector capable of distinguishing real images from CNN-generated ones without being dependent on a particular network architecture or training dataset.
-De-Fake [46] extends the detection pipeline by incorporating both the image content and generated prompts using CLIP to train a hybrid classifier.
-DIRE [54] and ZeroFake [47] exploit intrinsic differences between real and fake images revealed during the diffusion model reconstruction process to build detection models with improved generalization.
-PatchCraft [59] focuses on local texture patches, identifying subtle artifacts left by generative models in fine-grained regions.
-Recent works utilize VLMs as fake image detectors, such as AntifakePrompt [18], the method proposed by Jia et al.
-[29], and the approach by Ji et al.
-[28].
-AIGI-Holmes [60] proposes a complete framework to train a VLM for explainable and generalizable detection, aiming to produce human-verifiable justifications.
-The work by Yu et al.
-[57] develops a framework to enhance generalization and explainability by using a knowledge-guided detector and a forgery-aware prompt learner.
-FakeBench [33] and DFBench [51] introduced a large-scale benchmark to rigorously test the detection performance of LMMs against a wide range of modern generative models.
-Despite these progresses, existing methods exhibit inherent limitations.
-These approaches predominantly rely on internal, pixel-level visual features, while overlooking complementary external sources of evidence that can be critical for reliable forensic analysis.
-In contrast to standalone detector architectures, our work adopts an agent-based framework that emulates the investigative workflow of human forensic experts.
-The proposed AI agent leverages a diverse set of external tools to systematically collect, integrate, and reason over different evidence, which allows for a more robust, context-aware, and explainable framework for AI-generated image detection.
-LLM-Based Multi-Agent Frameworks
-Recent advances in large language models (LLMs) have catalyzed the development of agentic frameworks that simulate complex human-like workflows by coordinating multiple specialized agents.
-ReAct [56] introduces a framework that switches between reasoning and acting within language models.
-Other works such as MetaGPT [24] simulate various roles in a software company and build a multi-agent software development framework.
-In the security domain, several studies have developed multi-agent frameworks tailored to diverse security-related downstream tasks, including adversarial defense [58], harmful content detection [36], bias identification in generative models [52], fake news verification [32], and comparative evaluation of misinformation detection strategies [27].
-However, existing approaches do not specifically target zero-shot detection tasks nor address the unique challenges of AI-generated image detection.
-To the best of our knowledge, our work is the first to develop a training-free, zero-shot multi-agent framework specifically designed for AI-generated image detection.
-Combining the strengths of existing image detection methods and LLM agents, we present a novel and practical solution for addressing the evolving challenges in AI-generated image forensics.
 Conclusion
-In this work, we introduce AIFo, a novel multi-agent framework that advances AI-generated image detection by emulating human forensic reasoning.
-Unlike conventional methods, AIFo integrates diverse forensic tools and uses multi-agent collaboration to synthesize evidence.
-Our evaluation on a comprehensive benchmark spanning both lab and real-world settings shows that AIFo achieves 97.05% accuracy, substantially outperforming traditional classifiers and state-of-the-art VLMs like GPT-4o.
-The key contributions include a training-free, agent-based paradigm that ensures generalizability through procedural reasoning, a finding that multi-source evidence integration and structured debate mechanisms significantly enhance both accuracy and interpretability, and the establishment of a comprehensive benchmark that enables evaluation under real-world conditions.
-While our framework represents a significant advancement in AI-generated image detection, it also faces some limitations including computational overhead and dependency on external services.
-As generative AI technologies continue to evolve rapidly, the need for robust, interpretable, and adaptable detection systems becomes increasingly critical.
-Our agent-based approach offers a sustainable foundation for addressing these challenges, providing a framework that can generalize alongside advancing generative models while maintaining the transparency and reliability essential for real-world deployment in security-critical applications.
+This work presents AIFo, a multi-agent procedural reasoning framework that takes AI-generated image detection as an evidence-driven forensic decision process.
+Instead of proposing yet another detector, we show that explicitly modeling how forensic evidence is collected, evaluated, and reconciled leads to more reliable and interpretable decisions.
+Extensive experiments demonstrate that integrating heterogeneous evidence within a structured reasoning pipeline enables effective conflict resolution and decision revision.
+Our results suggest that robustness in AI-generated image detection is better achieved through multi-stage reasoning over evidence than through increasingly complex detection models alone.
+We believe that procedural, agent-based reasoning provides a promising foundation for developing forensic systems that remain effective as generative models continue to evolve.
+Limitation.
+The proposed framework has higher computational cost than single-model baselines due to multi-stage reasoning and external tool invocation.
+While our analysis shows that conflict resolution reasoning is only required for a subset of inputs, further optimization will be important for large-scale deployment.
+In addition, parts of the reasoning process, including the assessment of evidence sufficiency and consistency, rely on qualitative judgments produced by LLMs and may introduce variability across models or prompts.
+Exploring hybrid approaches that combine procedural reasoning with more formal uncertainty estimation and robustness guarantees is a promising avenue for future research.
 Ethical Considerations
@@
 - A real image refers to images created by humans, including photographs captured by cameras, photos that have been edited with software such as Photoshop, or human artistic creations such as hand-drawn sketches and paintings.
-- An AI-generated image refers to images that is fully or partially generated by AI models.
+- An AI-generated image refers to images that are fully or partially generated by AI models.
 Available Tools:
@@
 - A real image refers to images created by humans, including photographs captured by cameras, photos that have been edited with software such as Photoshop, or human artistic creations such as hand-drawn sketches and paintings.
-- An AI-generated image refers to images that is fully or partially generated by AI models.
+- An AI-generated image refers to images that are fully or partially generated by AI models.
 Please make a final judgment based on the following evidence collected from multiple tools:
@@
 - A real image refers to images created by humans, including photographs captured by cameras, photos that have been edited with software such as Photoshop, or human artistic creations such as hand-drawn sketches and paintings.
-- An AI-generated image refers to images that is fully or partially generated by AI models.
+- An AI-generated image refers to images that are fully or partially generated by AI models.
 Raw Evidence from tools: tool_results Arguments for ’AI-generated’:
@@
 MPF: Multi-picture format Metadata for multi-frame images.
-Appendix C Detailed Accuracy performance
-Table 17 provides a detailed breakdown of the accuracy performance of various methods across different image sources, including both in-the-lab and in-the-wild scenarios.
-Table 17: Detailed Accuracy performance of different methods on each image sources.
-In-the-Lab In-the-Wild
-Real Images AI Images Real Images AI Images
-Method Flickr30k ImageNet DIV2k GenImage FakeBench Holopix50k Flickr W. Commons Lexica Nightcafe Civitai
-CNNSpot [53] 0.9980 0.9980 1.0000 0.0475 0.1857 1.0000 1.0000 0.9980 0.0000 0.0020 0.0000
-PatchCraft [59] 0.9900 0.9140 0.7680 0.8525 0.5986 0.8600 0.9040 0.5960 0.0140 0.2360 0.3360
-DE-FAKE [46] 0.8980 0.7280 0.3480 0.7225 0.6443 0.7180 0.5700 0.4920 0.9840 0.9380 0.8360
-GPT-4.1 [8] 1.0000 0.9920 0.9880 0.8350 0.9086 0.9980 0.9860 0.9920 0.7840 0.9900 0.9500
-GPT-4o [39] 1.0000 0.9860 0.9960 0.8850 0.9471 0.9980 0.9820 0.9940 0.7520 0.9900 0.9420
-AIFo (ours) 1.0000 0.9840 0.9920 0.9475 0.9657 0.9960 0.9880 0.9940 0.8420 0.9880 0.9940
+Appendix C List of Selected Models from Hugging Face
+We select the top five most downloaded classification models for AI-generated image detection available on Hugging Face:
+  • haywoodsloan/ai-image-detector-deploy
+  • Organika/sdxl-detector
+  • legekka/AI-Anime-Image-Detector-ViT
+  • Smogy/SMOGY-Ai-images-detector
+  • NYUAD-ComNets/NYUAD_AI-generated_images_detector
+All models are publicly available on the Hugging Face Hub.
 Appendix D AI Model Sources
@@
 In-the-Lab AI Image Sources
-GenImage [61] BigGAN, GLIDE, VQDM, ADM, Midjourney, Wukong, and Stable Diffusion (v1.4, v1.5).
-FakeBench [33] ProGAN, StyleGANs, CogView2, FuseDream, VQDM, GLIDE, Midjourney, Stable Diffusion, DALL·E 2, and DALL·E 3.
+GenImage [ZCYHLLTHHW23] BigGAN, GLIDE, VQDM, ADM, Midjourney, Wukong, and Stable Diffusion (v1.4, v1.5).
+FakeBench [LLWLWRL24] ProGAN, StyleGANs, CogView2, FuseDream, VQDM, GLIDE, Midjourney, Stable Diffusion, DALL·E 2, and DALL·E 3.
 In-the-Wild AI Image Sources
-Lexica [9] Lexica Aperture Series (v3.5, v4, v5, Max).
-Nightcafe [10] DALL·E 2, DALL·E 3, Stable Diffusion, and various other community fine-tuned models.
-Civitai [5] A vast collection of community fine-tuned models, predominantly based on Stable Diffusion (including SDXL variants) series.
+Lexica Lexica Aperture Series (v3.5, v4, v5, Max).
+Nightcafe DALL·E 2, DALL·E 3, Stable Diffusion, and various other community fine-tuned models.
+Civitai A vast collection of community fine-tuned models, predominantly based on Stable Diffusion (including SDXL variants) series.
+Appendix E Detailed Accuracy Performance
+Table 17 provides a detailed breakdown of the accuracy performance of various methods across different image sources, including both in-the-lab and in-the-wild scenarios.
+Table 17: Detailed Accuracy performance of different methods on each image source.
+In-the-Lab In-the-Wild
+Real Images AI Images Real Images AI Images
+Method Flickr30k ImageNet DIV2K GenImage FakeBench Holopix50k Flickr W. Commons Lexica Nightcafe Civitai
+CNNSpot [WWZOE20] 0.9980 0.9980 1.0000 0.0475 0.1857 1.0000 1.0000 0.9980 0.0000 0.0020 0.0000
+PatchCraft [ZXLQZ24] 0.9900 0.9140 0.7680 0.8525 0.5986 0.8600 0.9040 0.5960 0.0140 0.2360 0.3360
+DE-FAKE [SLYZ22] 0.8980 0.7280 0.3480 0.7225 0.6443 0.7180 0.5700 0.4920 0.9840 0.9380 0.8360
+GPT-4o [GPT4o] 1.0000 0.9860 0.9960 0.8850 0.9471 0.9980 0.9820 0.9940 0.7520 0.9900 0.9420
+AIFo (ours) 1.0000 0.9840 0.9920 0.9475 0.9657 0.9960 0.9880 0.9940 0.8420 0.9880 0.9940